### v1

# SPIRIT: Short-term Prediction of solar IRradIance for zero-shot Transfer learning using Foundation Models

Abstract.
Traditional solar forecasting models are based on several years of site-specific historical irradiance data, often spanning five or more years, which are unavailable for newer photovoltaic farms. As renewable energy is highly intermittent, building accurate solar irradiance forecasting systems is essential for efficient grid management and enabling the ongoing proliferation of solar energy, which is crucial to achieve the United Nations’ net zero goals. In this work, we propose SPIRIT, a novel approach leveraging foundation models for solar irradiance forecasting, making it applicable to newer solar installations. Our approach outperforms state-of-the-art models in zero-shot transfer learning by about 70%, enabling effective performance at new locations without relying on any historical data. Further improvements in performance are achieved through fine-tuning, as more location-specific data becomes available. These findings are supported by statistical significance, further validating our approach. SPIRIT represents a pivotal step towards rapid, scalable, and adaptable solar forecasting solutions, advancing the integration of renewable energy into global power systems.
Solar Forecasting, Renewable Energy, Foundation Models, Transfer Learning, Zero-shot Learning, Fine-tuning, Deep Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Computing methodologies Machine learningccs: Applied computing Forecastingccs: Hardware Renewable energyccs: Computing methodologies Foundation modelsccs: Computing methodologies Transfer learning
Figure 1. Illustration of our system: A vision encoder (top-left) extracts embeddings from a sky camera image sampled from a diverse set spanning multiple locations and setups. Physics-inspired features are derived and integrated with auxiliary values, then merged with the image embedding (top-middle) into a unified representation. For nowcasting (right), a regressor predicts Global Horizontal Irradiance from this feature vector. For forecasting (bottom), a time-series model processes past feature vectors to create a context embedding, which is concatenated with a future covariate vector—constructed from known future values—to form the final latent representation. A regressor then maps this representation to future GHI values (bottom-right).
1. Introduction
The proliferation of solar energy is paramount for electrification and the global energy transition to meet the Net Zero commitments of the United Nations (Sadhukhan, 2022). As the world moves toward renewable sources, solar energy is notable for its accessibility and potential to significantly reduce carbon emissions (Sen, 2008). Expanding the solar energy infrastructure is crucial to mitigate the effects of climate change (Bashir et al., 2021) and meet the energy demands arising from sectors such as data centers (Agarwal et al., 2021), transportation (Lee et al., 2016), and buildings (Iyengar et al., 2017).
Unlike conventional power sources such as thermal and nuclear, solar energy has inherent shortcomings. Its intermittency, due to the daily and seasonal variations in sunlight, poses significant challenges for energy grid stability (Abido et al., 2022). One notable issue arising from the higher penetration of solar power is the “duck curve” (Iyengar et al., 2016), where the mismatch between solar energy production and peak energy demand leads to significant challenges in grid management. Although storage capacity is increasing, electricity grids typically operate as a just-in-time system where energy supply and demand must be balanced (Joskow, 2012). To ensure grid efficiency, renewable operators must pay a deviation penalty to discourage unplanned energy contributions, thereby maintaining a balanced and predictable energy supply (Yang et al., 2020). Thus, accurate short-term solar predictions are crucial for the efficient operation of the energy grid (Iyengar et al., 2014).
Existing approaches for short-term forecasting use Sky Cameras — i.e., a fish-eye lens camera positioned to look directly towards the zenith — which require extensive site-specific data to train models (Hammond et al., 2024; Gao and Liu, 2022). These models have demonstrated the ability to develop high-accuracy models, albeit using training data spanning multiple years. With the overall solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030 (isa, 2023), 90% of the solar farms worldwide will have negligible data to train custom models from scratch. Thus, lack of sufficient site-specific solar data underscores the need for approaches that do not compromise model performance.
With the advent of vision foundation models, we have seen improvement in accuracy of various Computer Vision tasks — such as feature extraction, object detection, etc. — using zero-shot and few-shot approaches (i.e., with limited or no custom training data) (Dosovitskiy, 2020; Zohar et al., 2023; Jeeveswaran et al., 2022). In addition, physics-inspired feature engineering has significantly improved model performance by incorporating domain-specific knowledge, leading to more accurate and interpretable predictions in real-world problems (Ompusunggu and Hostens, 2021; Erdmann et al., 2020). In this work, our hypothesis is as follows: Can we leverage state-of-the-art vision foundation models and physics-inspired features, along with transfer learning strategies, to reduce the dependence on site-specific sky camera imagery data?
To address these challenges, we introduce SPIRIT, a novel approach to solar irradiance forecasting with an inductive bias toward enhanced generalizability. In designing, implementing, and evaluating our approach, we make the following contributions:
(1) We develop a novel system that leverages foundation models and physics-informed features, eliminating the need for site-specific model training while enabling effective adaptation across diverse transfer learning scenarios. The flexibility of our framework ensures seamless integration of future advancements in vision models without requiring significant architectural modifications.
(2) Motivated by real-world deployment constraints, we demonstrate that SPIRIT can rapidly scale to new solar plant locations without prior sky camera data, significantly accelerating integration into operational workflows.
2. Related Work
Traditional methods for solar forecasting have relied heavily on Numerical Weather Prediction (NWP) models and satellite imagery (Markovics and Mayer, 2022). While these methods provide valuable insights, they often lack the spatial and temporal resolution required for accurate short-term forecasts. For instance, NWP models typically operate on a grid scale of several kilometers and update every few hours which may not capture rapid changes in cloud cover that affect solar irradiance (Kostylev et al., 2011). Over the past few years, several time series forecasting approaches have been used for solar forecasting. However, they typically operate on a time frame of multiple hours to day-ahead and are not suitable for capturing short-term variations in solar generation due to transient factors such as cloud cover (Iyengar et al., 2014; Falope et al., 2024).
Use of sky camera imagers for short-term solar forecasting has garnered significant attention in recent years due to their potential to enhance the accuracy of solar power predictions (Hammond et al., 2024; Gao and Liu, 2022; Nie et al., 2024). Sky cameras, equipped with fish-eye lenses, capture wide-angle images of the sky, providing valuable data on cloud cover and movement, which are critical factors in solar irradiance forecasting (Dev et al., 2019). Recent advancements have focused on leveraging sky cameras to address the limitations of traditional approaches. Hammond et al. (Hammond et al., 2024) and Gao et al. (Gao and Liu, 2022) demonstrated the potential of sky cameras in developing high-accuracy models for short-term solar forecasting. These studies utilized extensive site-specific data collected over multiple years to train their models, achieving significant improvements in forecast accuracy compared to traditional methods.
Siddiqui et al.(Siddiqui et al., 2019) proposed a deep learning framework using sky-camera images and auxiliary meteorological data to predict solar irradiance. Their approach employs a convolutional neural network (CNN) with dilated convolutions, followed by an LSTM for temporal forecasting up to four hours ahead. By training on 10 years of data, they demonstrated that incorporating auxiliary data such as temperature, wind speed, and relative humidity enhances generalization and stability in predictions. Similarly, Gao et al. (Gao and Liu, 2022) introduced a transformer-based architecture that integrates a clear sky model to estimate the residual irradiance beyond clear-sky assumptions. Trained on 10 years of data, their model achieves improved forecasting accuracy compared to earlier CNN-LSTM-based methods. Both works underscore the importance of leveraging sky images and auxiliary data for precise solar nowcasting and forecasting. Despite their promise, sky camera-based approaches face challenges related to data availability. With the global solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030, a large majority of the solar farms worldwide will have negligible historical data to train custom models from scratch.
Building upon these challenges, it becomes evident that addressing the limited availability of site-specific data is critical for advancing solar forecasting. Although the use of sky cameras and auxiliary data has substantially improved short-term predictions, the scalability of these methods remains constrained by the dearth of historical data at many solar installations. In this context, transfer learning emerges as a promising solution, as it enables the leveraging of knowledge from pre-trained models and the adaptation of learned representations across different datasets and locations. Notably, previous work such as Nie et al(Nie et al., 2024) has demonstrated that training on a fusion of multiple datasets yields models that perform better on each individual dataset, thereby highlighting the potential benefits of cross-dataset knowledge transfer.
3. SPIRIT Design
3.1. Key Concepts and Problem Setup
Nowcasting refers to the prediction of solar power generation over very short time horizons, typically ranging from a few minutes to a few hours (Lee et al., 2017). In contrast, short-term forecasting extends the prediction horizon to cover periods from one hour to 24 hours (Remund and Müller, 2012). Methods developed to provide forecasts utilize various data sources, such as satellite data (Lopes et al., 2021; Lee et al., 2017), weather station observations (Lee et al., 2017), and sky camera images (Gao and Liu, 2022; Xu et al., 2015; Siddiqui et al., 2019). Nowcasting and short-term forecasting are indispensable for managing the intermittency of solar power, allowing grid operators to perform better scheduling, dispatching, and balancing of energy resources (Dairi et al., 2020; Aouidad and Bouhelal, 2024).
Sky Camera: Sky cameras enhance nowcasting and short-term forecasting by capturing sky images with fish-eye lenses, providing detailed cloud movement, and sun position data. These images enable algorithms to track cloud dynamics and predict their trajectories, essential for estimating solar irradiance (Saraswat et al., 2023; Dev et al., 2019). Offering a low-latency alternative to weather satellites, sky cameras facilitate real-time monitoring. However, variations in camera setup and quality affect image appearance, as shown in Figure 4 in Appendix A.1. As a key tool in solar forecasting, sky cameras contribute to more reliable energy predictions (Rajagukguk et al., 2021). Further details are provided in Appendix B.
Irradiance measurements: Understanding solar irradiance requires distinguishing between three key measurements:
(1) Direct Normal Irradiance (DNI): The amount of solar radiation received per unit area on a surface perpendicular to the sun’s rays without being scattered or diffused by the atmosphere.
(2) Diffuse Horizontal Irradiance (DHI): The portion of solar radiation that reaches a horizontal surface after being scattered by molecules, aerosols, and clouds in the atmosphere. Unlike DNI, DHI comes from all directions in the sky and plays a crucial role during overcast conditions when direct sunlight is obstructed.
(3) Global Horizontal Irradiance (GHI): The total solar radiation received on a horizontal surface, combining both direct and diffuse components. GHI is the sum of DNI, projected onto a horizontal plane, and DHI:
(1) $GHI=DNI\times\cos(\theta)+DHI$
where $\theta$ is the angle between the direction of incoming solar radiation and the vertical, called the zenith angle.
GHI is the most commonly used irradiance measure in solar energy applications, as it directly influences photovoltaic (PV) panel performance and solar power generation, making it the primary focus of research in irradiance forecasting. Henceforth, unless explicitly stated otherwise, any mention of irradiance or solar irradiance refers specifically to Global Horizontal Irradiance.
Photovoltaic Power Output: PV power output refers to the electricity generated by solar panels from incoming solar radiation. While it is primarily driven by GHI (Vilanova et al., 2020), factors like temperature, and system losses also play a role. Under stable conditions, the relationship between GHI and PV output is roughly linear (Razak et al., 2016; Natheer Tuaimah and Al-Saidi, 2019). Since PV output is a more actionable metric for grid management and energy planning, predicting it directly is often more desirable.
3.2. Nowcasting Architecture
We propose an architecture that encodes sky images into vector representations, which are augmented with auxiliary data and physics-based features. This representation captures information about the GHI, which is then effectively extracted by a regression model.
Let $\mathcal{X}$ be the set of sky camera images, and $\mathcal{D}$ be the dataset, defined as $\mathcal{D}=\{(X_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{N}$, where $X_{i}\in\mathcal{X}$ represents the $i$-th sky image, $\mathbf{A}_{i}\in\mathbb{R}^{k}$ corresponds to the auxiliary features such as azimuth and zenith angles of the Sun, and $y_{i}\in\mathbb{R}^{+}$ are the corresponding solar irradiance measurements.
An encoder function $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$ is defined that assigns a $d$-dimensional embedding vector to each image $X\in\mathcal{X}$:
$\mathbf{Z}=E(X),\quad\mathbf{Z}\in\mathbb{R}^{d}$
To leverage domain knowledge in solar power prediction, we introduce a set of additional features, $\mathbf{P}$, derived from the auxiliary measurements $\mathbf{A}$. These features incorporate established solar engineering principles, such as clear sky irradiance, and panel tilt and orientation, as defined in Subsection 3.4. The feature vector $\mathbf{P}$ is given by:
$\mathbf{P}\in\mathbb{R}^{p}$
where $p$ represents the number of physics-based features extracted from the auxiliary data.
The final feature representation $\mathbf{f}\in\mathbb{R}^{d+k+p}$ is constructed by concatenating the image embedding $\mathbf{Z}$, raw auxiliary measurements $\mathbf{A}$, and the physics-based features $\mathbf{P}$:
$\mathbf{f}=\mathbf{Z}\oplus\mathbf{A}\oplus\mathbf{P}$
where $\oplus$ denotes the concatenation operation. This combined representation leverages data-driven features, visual features, and domain-specific engineering knowledge, providing a comprehensive characterization of each sample $(X_{i},\mathbf{A}_{i},y_{i})\in\mathcal{D}$.
A regression function $R_{\omega}:\mathbb{R}^{d+k+p}\rightarrow\mathbb{R}^{+}$, parameterized by weights $\omega$, is defined such that:
$\hat{y}=R_{\omega}(\mathbf{f})=R_{\omega}(E(X)\oplus\mathbf{A}\oplus\mathbf{P})$
Nowcasting loss function $\mathcal{L}_{nowcast}(\omega)$ is defined as the average of the individual regression losses for each sample, where each individual loss measures the discrepancy between the predicted $\hat{y}_{i}=R_{\omega}(\mathbf{f}_{i})$ and the true value $y_{i}$:
$\mathcal{L}_{nowcast}(\omega)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(R_{\omega}(\mathbf{f}_{i}),y_{i})$
where $\mathcal{L}(R_{\omega}(\mathbf{f}_{i}),y_{i})$ is the regression loss for the $i$-th sample. To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{nowcast}(\omega)$ using gradient-based methods.
3.3. Forecasting Architecture
Our forecasting architecture processes sequences of sky images to predict GHI across multiple future intervals. Each image is encoded using the embedding and augmentation approach from Section 3.2. A time-series model captures a latent representation of past features, while predictable future covariates, such as the zenith angle, are precomputed and integrated as a vector. The combined past and future representations are then input into a regressor to generate GHI predictions.
A sequence of $T$ images $X_{1:T}=\{X_{1},X_{2},\dots,X_{T}\}$ along with their corresponding auxiliary features $\mathbf{A}_{1:T}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T}\}$, where each $\mathbf{A}_{t}\in\mathbb{R}^{k}$ represents the auxiliary feature vector at time $t$, is given.
An encoder function $E$ generates the vector representation $\mathbf{Z}_{t}=E(X_{t})\in\mathbb{R}^{d}$ for each image at time step $t=1,2,\dots,T$. The physics-based features $\mathbf{P}_{t}$ are derived from auxiliary measurements $\mathbf{A}_{t}$. The final feature vectors $\mathbf{f}_{t}\in\mathbb{R}^{d+k+p}$ are obtained by concatenating the image embedding, auxiliary data, and physics-based features:
$\mathbf{f}_{t}=\mathbf{Z}_{t}\oplus\mathbf{A}_{t}\oplus\mathbf{P}_{t}$
where $\oplus$ denotes concatenation, providing a comprehensive characterization of each sample $(X_{t},\mathbf{A}_{t},y_{t})\in\mathcal{D}$.
Thus, the collection of feature vectors over the sequence of $T$ time steps is given by:
$\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$
where $\mathbf{F_{1:T}}$ represents the set of concatenated feature representations created for each timestamp in the sequence.
Given the collection of feature vectors $\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$, a time-series model $\mathcal{M}$ is used to encode the observed sequence into a latent vector $\mathbf{L}\in\mathbb{R}^{m}$, which captures the full context of the input data series while retaining its temporal patterns and dependencies:
$\mathbf{L}=\mathcal{M}(\mathbf{F_{1:T}})\in\mathbb{R}^{m}$
where $\mathcal{M}$ represents the time-series model that transforms the observed sequence of feature vectors $\mathbf{F_{1:T}}$ into a compact representation in the latent space $\mathbb{R}^{m}$.
To integrate known future information, derived from the spatiotemporal context of time and location, future covariate vectors $\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$ are constructed for each forecast time $T+\tau_{i}$, . The full covariate vector $\mathbf{C}\in\mathbb{R}^{q\cdot H}$ is then formed by concatenating these individual representations across all forecast horizons:
$\mathbf{C}=\bigoplus_{i=1}^{H}\mathbf{C}_{T+\tau_{i}},\quad\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$
We concatenate the future covariate vector $\mathbf{C}$ with the latent representation of the past time steps $\mathbf{L}$, forming the final vector that encompasses all relevant information:
$\mathbf{h}=\mathbf{L}\oplus\mathbf{C}$
This ensures that both past contextual information as well as known future data contribute to the forecasting process.
Next, a regression function $R_{\omega}:\mathbb{R}^{m+q\cdot H}\to\mathbb{R}^{H}$, parameterized by $\omega$, is applied to the vector $\mathbf{h}\in\mathbb{R}^{m+q\cdot H}$ to generate the corresponding predicted GHI values. The regressor outputs a vector $\hat{\mathbf{y}}\in\mathbb{R}^{H}$ of predicted GHI values for the forecast time intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$:
$\hat{\mathbf{y}}=R_{\omega}(\mathbf{h})=\left[\hat{y}_{T+\tau_{1}},\hat{y}_{T+\tau_{2}},\dots,\hat{y}_{T+\tau_{H}}\right]\in\mathbb{R}^{H}$
where each $\hat{y}_{i}$ corresponds to the irradiance forecast for the time interval $T+\tau_{i}$, and the vector $\hat{\mathbf{y}}$ represents the full set of predicted irradiance values across all forecast intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$.
Forecasting loss function $\mathcal{L}_{forecast}(\omega)$ is defined as the mean of the individual regression losses computed over all forecast intervals $T+\tau_{j}$ for each sample $i$ Specifically, the total loss is given by:
$\mathcal{L}_{forecast}(\omega)=\frac{1}{N\cdot H}\sum_{i=1}^{N}\sum_{j=1}^{H}\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$
where $\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$ is the individual regression loss for the forecast interval $T+\tau_{j}$ for sample $i$ To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{forecast}(\omega)$ using gradient-based optimization methods. The complete architecture is illustrated in Figure 1.
The Significance of Generalized Encoders: A key distinction of our approach is that in prior work (Gao and Liu, 2022; Hasan, 2023; Siddiqui et al., 2019), the encoder $E$ is a vision model typically trained on data from a specific location and camera setup. Furthermore, studies aiming for generalizability typically rely on training models using a fusion of solar datasets from multiple locations (Nie et al., 2024; Despotovic et al., 2024). In contrast, we argue, and later demonstrate, that leveraging a foundation model, a highly generalizable feature extractor, provides a more robust $E$ function. A foundation model not only matches the performance of site-specific encoders at a given location with a particular setup but also demonstrates an unparalleled advantage in generalizing across diverse locations and camera setups.
3.4. Physics-inspired Feature Engineering
Clear sky models (Ineichen and Perez, 2002; Stein et al., 2012; Perez et al., 2002; Mueller et al., 2004) are mathematical models that estimate the theoretical solar irradiance at a given location under cloud-free conditions, serving as a representation of the maximum possible radiation reaching the Earth’s surface. These models leverage fundamental atmospheric physics and employ mathematical formulations based on solar geometry (Stein et al., 2012), atmospheric transmittance (Stein et al., 2012), and radiative transfer (Stein et al., 2012) to derive estimations of GHI, DNI and DHI under clear sky conditions. The Ineichen clear sky model (Ineichen and Perez, 2002) requires inputs such as latitude, longitude, time, and date, which are readily available. This allows clear sky irradiance values to be readily computed and incorporated into our model as features, providing a reference for expected irradiance levels in the absence of cloud interference.
Physics behind solar irradiance: Solar irradiance, the power per unit area received from the Sun in the form of electromagnetic radiation, is measured in watts per square meter $(W/m^{2})$. The amount of solar irradiance received by a solar panel depends on additional site-specific factors, including the panel’s tilt and orientation angle, the Sun’s altitude and azimuth, and the geographic location’s latitude and longitude. We first look at the angle of incidence ($\theta$(Laboratories, [n. d.]), i.e. the angle between the incoming solar rays and the normal to the surface of the solar panel. It can be calculated using the following formula:
(2) $cos(\theta)=cos(\theta_{z})\cdot cos(\beta)+sin(\theta_{z})\cdot sin(\beta)\cdot cos(\gamma-\alpha)$
where $\theta_{z}$ and $\gamma$ are the solar zenith and azimuth angles respectively. While $\beta$ and $\alpha$ are the tilt and azimuth angles of the panel.
We calculate the effective irradiance by adding the three main components: direct, diffuse, and reflected irradiance (see below):
(3) $I_{panel}=DNI\cdot cos(\theta)+DHI\cdot\frac{1+cos(\beta)}{2}+GHI\cdot\rho\cdot\frac{1-cos(\beta)}{2}$
where $I_{panel}$ is the effective irradiance and $\rho$ is the ground reflectance.
4. SPIRIT Implementation
4.1. Nowcasting
In our approach, we utilize the pre-trained Google Vision Transformer (ViT) (Dosovitskiy, 2020), a model with 632 million parameters, to generate embeddings for sky camera images. To reduce sensor dependence and focus on image features, we exclude meteorological sensor data, incorporating only auxiliary variables such as zenith and azimuth angles, clear sky irradiance, panel tilt, and orientation. These image embeddings are subsequently concatenated with the auxiliary vector to form the final feature representation. The combined feature vectors, paired with their corresponding ground truth GHI values, are then used to train an XGBoost regressor within a supervised learning framework. The model is optimized by minimizing the Mean Squared Error (MSE) loss function, which measures the difference between the predicted and actual GHI values.
4.2. Forecasting
For forecasting, we employ the Google Vision Transformer (ViT) (Dosovitskiy, 2020) to generate image embeddings, which are subsequently concatenated with the auxiliary variables to form a comprehensive feature representation. To account for temporal dependencies, we input a sequence of six images, representing a 1-hour context window, into a transformer-based time-series encoder (Vaswani, 2017). This encoder processes the temporal sequence and learns a latent representation of the past context, which is then fused with a future covariate vector that includes azimuth and zenith angles, as well as clear sky GHI. The resulting representation is passed through a multi-layer perceptron (MLP) to predict the solar irradiance for the 1-hour, 2-hour, 3-hour, and 4-hour forecast intervals. This implementation exemplifies one approach in our framework, with additional variations incorporating different vision encoders of varied sizes in the ablation studies detailed in Section 7.
5. Evaluation Methodology
5.1. Datasets
We evaluate our methods using three publicly available datasets: TSI880 (Andreas and Stoffel, 1981), ASI16 (Andreas and Stoffel, 1981), and SKIPP’D (Nie et al., 2023). The TSI880 and ASI16 datasets, both collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado, provide sky images captured every 10 minutes along with corresponding GHI values and auxiliary data such as air temperature and relative humidity and only differ in camera setup and sensors, with the ASI16 dataset capturing higher-resolution images. The SKIPP’D dataset, collected from Stanford University, consists of raw sky images captured every minute and PV power output data, prioritizing finer temporal granularity at the expense of image quality. For more details, refer to Appendix A.
We utilize the TSI880 and ASI16 datasets to investigate the impact of camera setup at the same location. To explore location and task shifts, we use the SKIPP’D dataset to evaluate the performance of models trained on GHI data in predicting PV power output. The SKIPP’D dataset features lower-resolution images and lacks meteorological data, thereby presenting a more challenging task by limiting the contextual information typically leveraged by prior models (Gao and Liu, 2022; Siddiqui et al., 2019). To ensure the models learn from higher-quality, information-rich datasets, we train exclusively on the TSI and ASI datasets while evaluation is done across all the datasets, including the more challenging SKIPP’D, allowing us to assess how well the models generalize to lower-quality data and increased domain shifts.
5.2. Performance Metrics
We assess the effectiveness of the predicted values using the normalized Mean Absolute Percentage error (nMAP), defined as:
(4) $\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$
where $y_{i}$ represents the actual value and $\hat{y}_{i}$ represents the predicted value for the $i$-th sample, with $i\in\{1,\dots,N\}$. It is commonly used for solar irradiance prediction as the normalization ensures that models can be assessed uniformly on datasets with varied value ranges, avoiding biased assessments due to scale differences.
5.3. Baselines
To benchmark our proposed method, we compare its performance against the state-of-the-art baseline, Gao et al. (Gao and Liu, 2022), who demonstrate state-of-the-art performance for nowcasting and forecasting by training a vision transformer (Dosovitskiy, 2020) from the ground up using 10 years of site-specific data (Andreas and Stoffel, 1981; Gao and Liu, 2022; Siddiqui et al., 2019). Their approach for forecasting utilizes a temporal transformer (Vaswani, 2017), also trained on the same duration of data. To ensure a fair comparison, we reproduced their architecture and conducted experiments under the same conditions for both Gao et al.’s (Gao and Liu, 2022) model and SPIRIT.
Table 1. Nowcasting performance across multiple datasets: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset for a year are evaluated with nMAP both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017. We observe comparable performance when tested in the training setup, but our model demonstrates significantly better zero-shot performance in a new location.
Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
TSI ASI 27.17 (-62.49) 89.66
SKIPP’D 35.94 (-60.94) 96.43
TSI 9.04 (+0.08) 8.96
ASI TSI 28.86 (-46.65) 75.51
SKIPP’D 32.98 (-57.69) 90.67
ASI 9.08 (+0.95) 8.13
Table 2. Forecasting performance across multiple datasets and forecast intervals: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset are evaluated with nMAP error both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017 across four forecast intervals: 1hr, 2hr, 3hr, and 4hr.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 29.99(-5.75) 35.74
SKIPP’D 32.93(-5.95) 38.88
TSI 18.96(-1.00) 19.96
ASI TSI 26.85(-2.19) 29.04
SKIPP’D 27.33(-14.35) 41.68
ASI 19.23 (+0.02) 19.21
2hr TSI ASI 31.71(-5.89) 37.60
SKIPP’D 29.01(-14.80) 43.81
TSI 21.77(-0.87) 22.64
ASI TSI 28.64(-1.01) 30.65
SKIPP’D 26.29(-21.63) 47.92
ASI 21.51(-0.47) 21.98
3hr TSI ASI 34.41(-3.36) 37.77
SKIPP’D 30.26(-17.10) 47.36
TSI 25.46(-0.84) 26.30
ASI TSI 31.65(-1.5) 33.15
SKIPP’D 30.26(-22.89) 53.15
ASI 24.78(-0.89) 25.67
4hr TSI ASI 38.00(-1.58) 39.58
SKIPP’D 34.63(-17.15) 51.78
TSI 29.89(-1.69) 31.58
ASI TSI 35.86(-0.99) 36.85
SKIPP’D 36.97(-13.20) 50.17
ASI 29.29(-1.73) 31.02
5.4. Zero-shot Transfer Learning
To evaluate the zero-shot generalization performance of our models, we analyze two distinct transfer learning scenarios. The first scenario examines intra-location generalization, where the models are trained and tested in the same geographic location but under varying camera setups. While the environmental conditions remain consistent, variations in camera setup, viewing angles, and image resolutions exist between the training and testing phases. When image-based models are trained on data from a particular camera setup, they learn to associate specific regions of the image with key features—such as the position of the sun, cloud formations, or atmospheric conditions—that influence the predicted output. However, when the camera setup is altered, the spatial mapping of these features within the image shifts. To assess how well the models handle such variations, we train them using the TSI dataset and evaluate them on the ASI dataset, and vice versa.
The second scenario focuses on cross-location and cross-task generalization, where models trained in one geographic location are tested in another with different environmental and sensor characteristics. We train on the TSI and ASI datasets and evaluate on the SKIPP’D dataset, with the task shifting from predicting GHI to PV power output. Since GHI and PV output have a nearly linear correlation (Vilanova et al., 2020), this serves as a valid example of heterogeneous transfer learning. To account for the significant scale difference between GHI and PV output, model outputs are normalized for comparability. We conduct experiments for both nowcasting and forecasting tasks, training the models for one year and testing them on another year to account for seasonal variations, thus ensuring a fair evaluation. The nMAP errors are reported in Table 1 for nowcasting and Table 2 for forecasting, comparing SPIRIT with the state-of-the-art in both-the zero-shot transfer learning setups and the traditional settings, where the models are trained and tested using data from the same location and setup, but different years.
Figure 2. We compare the nowcasting performance of SPIRIT and Gao et al. using nMAP error. The solid lines represent the average performance across different fine-tuning training sizes, measured in weeks of data. The shaded regions indicate the xconfidence interval, reflecting variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting fine-tuning data from different starting points throughout the year.
Figure 3. We compare the forecasting performance of SPIRIT and Gao et al. using nMAP error across different forecast intervals. Subfigures (a), (b), (c), and (d) correspond to 1-hour, 2-hour, 3-hour, and 4-hour forecasting, respectively. The solid lines represent the average performance for each forecast interval, with varying fine-tuning training sizes measured in weeks of data. The shaded regions denote the 95% confidence interval, illustrating the variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting weeks of contiguous fine-tuning data from different starting points throughout the year. SPIRIT exhibits consistently low variance compared to the baseline, particularly in settings with severely limited data, demonstrating its ability to maintain stability. In contrast, the baseline shows high variance, indicating uncertainty in its predictions.
5.5. Fine-tuning with Limited Data
Building upon our zero-shot transfer learning experiments, we now investigate the adaptability of our models in a fine-tuning framework, where a limited amount of labeled data from the target domain is available for fine-tuning. This scenario closely resembles practical deployment conditions, where prolonged data collection is often infeasible, and models must quickly adapt to new locations with minimal supervision. We evaluate transfer learning with limited data in two scenarios: intra-location adaptation and cross-location adaptation, as in Subsection 5.4.
For both experimental setups, we perform fine-tuning using progressively increasing amounts of labeled data from the target domain—specifically, one, two, three, and four weeks of data from a full year for nowcasting, and two, four, eight, twelve, and sixteen weeks of data for forecasting with testing done on the remaining data from the year. Given the greater complexity of forecasting, we extend the fine-tuning experiment to a larger time frame. Additionally, due to the requirement for temporal consistency in the time series data, as discussed in Subsection A.2, the number of nowcasting samples for a given time period exceeds that of forecasting samples. We implement a selective fine-tuning approach, where only the regressors (see Figure 1) are updated, while the rest of the model is frozen. This ensures that the pre-trained feature representations, which capture generalizable spatiotemporal patterns, remain intact while allowing the model to adapt to location- and camera-specific variations. As demonstrated in prior work (Nie et al., 2024; Zhou et al., 2020; Sarmas et al., 2022), fine-tuning only the final layers achieves competitive adaptation performance while mitigating the risk of overfitting to the limited target data The nowcasting metrics are shown in Figure 2, and the forecasting metrics are depicted in Figure 3.
6. Results
6.1. Zero-shot Transfer Learning
Tables 1 and 2 present the results for zero-shot transfer learning, demonstrating that our model consistently outperforms the state-of-the-art baseline across both cross-location and cross-setup scenarios in both nowcasting and forecasting tasks. When transitioning between camera setups within the same location, our model consistently shows better performance relative to the baseline. However, more notably, when moving across different locations, our model achieves up to 45% improvement. In this more complex cross-location setting, our model significantly outperforms the baseline, highlighting its superior generalizability and robustness. Furthermore, even in the traditional setup where models are trained and tested on the same location, our approach demonstrates enhanced forecasting performance, further emphasizing its effectiveness across diverse deployment conditions.
6.2. Fine-tuning with Limited Data
For the analysis of fine-tuning results, we merge cross-setup and cross-location scenarios to ensure a sufficient number of data points for robust confidence interval plots, as depicted in both the nowcasting (Figure 2) and forecasting (Figure 3) tasks. Since nowcasting is a relatively simpler task, both models exhibit rapid improvement within the first week. However, the baseline model reaches performance saturation early, at approximately 45%, while our model continues to reduce its error, achieving a significant improvement, dropping below 20% within four weeks.
In forecasting, SPIRIT consistently outperforms the baseline, demonstrating notably lower variance, particularly in data-limited settings (0-2 weeks of data). This underscores SPIRIT’s superior stability and reliability, with its nMAP error remaining consistently below that of the baseline. In contrast, the baseline model exhibits higher variance, indicating greater inconsistency and confusion in its performance when limited data is available. Both models experience a typical performance decline as the forecasting horizon extends from 1-hour to 4-hour forecasts, driven by the increased uncertainty over longer time horizons. Nonetheless, SPIRIT’s consistently lower variance and sustained performance highlight its robustness and its ability to adapt more effectively to challenging conditions. The transition from a zero-shot configuration to fine-tuning results in noticeable performance improvements; however, the gains diminish after approximately eight weeks of fine-tuning, suggesting that extended fine-tuning beyond this period yields only marginal additional benefits. All the results are in Appendix C.
7. Ablation Studies
7.1. Investigating Different Vision Encoders
We examine the impact of different vision models on SPIRIT’s performance, also highlighting the versatility of our system across different foundation models. We evaluate the CNN-based ResNet-152 (He et al., 2016), the vision transformer-based DINOv2 Giant (Oquab et al., 2023), and our implementation using Google ViT-Huge (Dosovitskiy, 2020). Results summarized in Table 3, demonstrate that the ViT-based models consistently outperform the ResNet-152 CNN model, which can be attributed to the superior capability of ViT architectures in capturing global image features (Jeeveswaran et al., 2022).
Table 3. We explore the impact of using different vision encoders on the overall model performance for nowcasting and forecasting, with training on TSI 2020 and testing on TSI 2021, measured by nMAP error.
Model Nowcast Forecast
+1hr +2hr +3hr +4hr
ResNet-152 10.50 24.56 27.82 31.23 35.85
DINOv2 Giant 9.74 21.22 23.56 27.93 33.13
Google ViT-Huge 9.32 19.96 22.64 26.30 31.58
7.2. Foundation Model Size
Table 4 presents an analysis of how the size of the foundation model influences the performance of our nowcasting and forecasting architectures. Although increasing model size has traditionally been linked to performance gains, we observe that beyond a certain threshold, further scaling yields diminishing returns. This suggests that larger models do not always lead to better performance. In fact, models with 304M and 86M parameters outperform their larger counterparts with 632M parameters in forecasting and nowcasting, respectively. This aligns with recent work, which highlights that adjusting model size based on a computational budget, rather than blindly increasing model size, can lead to more efficient architectures with reduced inference costs (Alabdulmohsin et al., 2023).
Table 4. We evaluate the impact of varying size of the Google ViT vision encoder on the overall performance of the model for both nowcasting and forecasting tasks, with training on TSI 2020 and testing performed on TSI 2021.
Model Parameters Nowcast Forecast
+1hr +2hr +3hr +4hr
86M 9.14 21.92 24.07 28.73 34.50
304M 9.45 19.58 21.95 25.54 30.60
632M 9.32 19.96 22.64 26.30 31.58
8. Conclusion
This work addresses a critical challenge in solar irradiance forecasting: adapting models to new geographic locations with no prior data. By utilizing transfer learning and pre-trained models, SPIRIT generalizes well to new locations, reducing the reliance on large, location-specific datasets. As more site-specific data becomes available post-deployment, the system can be effectively fine-tuned, improving prediction accuracy and supporting better energy yield estimates and operational planning. Additionally, SPIRIT’s modular design allows for the seamless integration of any emerging vision models, ensuring that the framework remains up-to-date with the latest advancements. This scalable solution for solar irradiance forecasting can accelerate the deployment of solar farms—particularly in remote and emerging markets. SPIRIT supports the transition to renewable energy by enhancing the reliability, cost-effectiveness, and accessibility of solar energy generation.
9. Future Work and Limitations
One key limitation is that the datasets used for evaluation are all from North America, largely due to the limited availability of publicly accessible datasets from other regions. Specifically, the solar movement patterns and dynamics change in the Southern Hemisphere and need to be studied. To improve the generalizability of our system, future work will incorporate data from other continents. Additionally, while our model performs well, the use of foundation models introduces real-time inference costs and computational overheads. Future efforts will focus on reducing computational efficiency, enabling deployment on resource-constrained edge devices without sacrificing accuracy.
Appendix A Dataset Details
Table 5. A Comparative Overview of the TSI880, ASI16, and SKIPP’D Datasets: Key Attributes Including Geographical Location, Data Provided, Image Resolution, Collection Frequency, and Annual Sample Size
Attribute TSI880 Dataset ASI16 Dataset SKIPP’D Dataset
Location Golden, Colorado, USA Golden, Colorado, USA Stanford, California, USA
Data Type Sky images & Irradiance data Sky images & Irradiance data Sky images & PV power output
Data Frequency 10-minutes 10-minutes 1-minute
Image Resolution 288x352 1536x1536 64x64
Camera Model Aero-Laser TSI-880 EKO ASI-16 Hikvision DS-2CD6362F-IV
Number of Samples / Year 24,948 25,107 121,125
A.1. Overview of Datasets
TSI880 Dataset: The TSI880 dataset is collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado. The camera captures an image every 10 minutes from 7:50 to 16:40 daily, providing raw sky images along with corresponding global horizontal irradiance values. Additionally, the dataset includes auxiliary information such as air temperature, relative humidity, azimuth angle, and zenith angle.
ASI16 Dataset: The ASI16 dataset is also sourced from the Solar Radiation Research Laboratory in Golden, Colorado, but it differs in that the camera setup captures images at a higher resolution. Similar to the TSI880 dataset, it provides global horizontal irradiance values and auxiliary data including azimuth angle, zenith angle, air temperature, relative humidity, and average wind speed.
SKIPP’D Dataset: The SKIPP’D dataset consists of raw sky images and photovoltaic (PV) power output data collected from Stanford University, California, USA. Images are captured every minute with a resolution of 64×64 pixels, emphasizing finer temporal granularity at the expense of lower image resolution.
Figure 4. Examples of sunny, partly cloudy, and overcast conditions, captured by different sky cameras, are shown from left to right, across the three datasets: TSI, ASI, and SKIPP’D, displayed from top to bottom.
A.2. Temporal Consistency in Forecasting
Valid samples for forecasting are formed such that all the data points from time steps $1$ to $T$, and their corresponding forecast intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$, fall within the same day. This is an essential requirement because the predictions for future intervals rely on the assumption that both historical and forecast data belong to the same day. Using data from the current day to predict values for the following day is not a valid forecasting approach, as the discontinuity between days renders such predictions unreliable. Any samples that violate this condition are considered invalid and are excluded from training or evaluation.
Appendix B Clear Sky Global Horizontal Irradiance
Clear Sky Global Horizontal Irradiance (GHI) is the solar irradiance received on a horizontal surface under cloud-free conditions. Most of the time, it serves as an upper bound for the actual GHI at a given location and time.
Clear Sky GHI plays a key role in solar forecasting by serving as a baseline for estimating how much clouds reduce solar irradiance. By comparing actual irradiance with Clear Sky GHI, we can get an estimate of the impact of cloud cover, which helps in enhancing short-term predictions, and improving the accuracy of forecasting models.
Given the latitude and longitude of a location, the clear sky values can be estimated for any timestamp. This becomes very useful in solar forecasting, as this value would give a reference of how much the prediction needs to be.
Clear Sky GHI is computed using mathematical models incorporating solar position, atmospheric transmittance, and radiative transfer principles. A common approach is the Ineichen-Perez model (Stein et al., 2012):
(5) $GHI_{\text{clear}}=I_{0}\cdot\tau\cdot\cos(\theta_{z})$
where $I_{0}$ is the extraterrestrial irradiance (W/m²), $\tau$ is the atmospheric transmittance factor, $\theta_{z}$ is the solar zenith angle.
B.1. Extraterrestrial Irradiance ($I_{0}$)
Extraterrestrial irradiance ($I_{0}$) is the solar irradiance just outside Earth’s atmosphere, slightly varying due to Earth’s elliptical orbit around the Sun. It is given by:
(6) $I_{0}=S_{c}\cdot\left(1+0.033\cos\left(\frac{2\pi n}{365}\right)\right)$
where $S_{c}=1367$ W/m² (solar constant), $n$ is the day of the year (1 for January 1, 365 for December 31).
B.2. Atmospheric Transmittance ($\tau$)
The atmospheric transmittance $\tau$ accounts for the attenuation of solar radiation by the atmosphere. It is often estimated using empirical models, such as the Ineichen-Perez model (Stein et al., 2012):
(7) $\tau=a\cdot e^{-b\cdot m}$
where $a,b$ are empirical coefficients dependent on location and aerosol content, $m$ is the air mass, given by (Ineichen and Perez, 2002):
(8) $m=\frac{1}{\cos(\theta_{z})+0.15(93.885-\theta_{z})^{-1.253}}$
where $\theta_{z}$ is the solar zenith angle.
Appendix C Fine-tuning Detailed Results
C.1. Nowcasting
To understand the impact of fine-tuning duration and the training size, we conducted a series of experiments by varying the amount of training data used for fine-tuning, by using subsets of the data consisting of 1, 2, 3, and 4 weeks.
Our results show that even with only one week of training data at a new location, the fine-tuned model performs remarkably well. Furthermore, in all experimental configurations, our model significantly outperforms the baseline.
Detailed results for these experiments are presented in Tables 6-9.
Table 6. Nowcasting Performance with 1 week training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
TSI ASI 20.23 52.01
SKIPP’D 29.89 63.82
ASI TSI 14.99 27.98
SKIPP’D 27.51 40.92
Table 7. Nowcasting Performance with 2 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
TSI ASI 18.96 51.45
SKIPP’D 29.07 62.91
ASI TSI 14.91 27.71
SKIPP’D 26.41 40.25
Table 8. Nowcasting Performance with 3 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
TSI ASI 16.52 50.38
SKIPP’D 27.42 62.05
ASI TSI 14.59 27.53
SKIPP’D 25.68 39.89
Table 9. Nowcasting Performance with 4 weeks training
Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
TSI ASI 15.63 50.01
SKIPP’D 26.51 61.17
ASI TSI 14.12 27.28
SKIPP’D 24.32 39.43
C.2. Forecasting
Table 10. Forecasting Performance with 2 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 31.15 33.86
SKIPP’D 32.35 38.24
ASI TSI 24.47 36.18
SKIPP’D 27.00 30.48
2hr TSI ASI 32.70 36.44
SKIPP’D 29.41 39.06
ASI TSI 25.93 36.71
SKIPP’D 25.96 33.55
3hr TSI ASI 34.41 38.24
SKIPP’D 31.53 39.84
ASI TSI 30.45 41.46
SKIPP’D 30.03 39.76
4hr TSI ASI 38.19 43.76
SKIPP’D 36.83 41.76
ASI TSI 36.44 45.89
SKIPP’D 36.84 44.16
Table 11. Forecasting Performance with 4 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 22.17 29.03
SKIPP’D 32.44 39.82
ASI TSI 27.65 35.77
SKIPP’D 26.54 30.70
2hr TSI ASI 25.13 32.69
SKIPP’D 29.56 40.21
ASI TSI 31.06 36.62
SKIPP’D 25.53 33.63
3hr TSI ASI 30.12 38.64
SKIPP’D 31.79 40.18
ASI TSI 34.47 38.76
SKIPP’D 29.73 39.70
4hr TSI ASI 36.14 41.92
SKIPP’D 37.24 41.31
ASI TSI 39.72 40.02
SKIPP’D 36.67 44.16
Table 12. Forecasting Performance with 8 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 22.62 32.45
SKIPP’D 33.56 36.94
ASI TSI 26.38 35.70
SKIPP’D 26.61 31.25
2hr TSI ASI 25.15 33.58
SKIPP’D 30.65 38.06
ASI TSI 26.68 35.26
SKIPP’D 25.30 33.95
3hr TSI ASI 28.66 35.57
SKIPP’D 32.64 39.29
ASI TSI 29.81 36.44
SKIPP’D 29.25 39.85
4hr TSI ASI 34.76 39.41
SKIPP’D 37.80 41.63
ASI TSI 34.97 38.23
SKIPP’D 36.23 44.25
Table 13. Forecasting Performance with 12 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 22.03 34.63
SKIPP’D 33.76 37.35
ASI TSI 24.87 35.24
SKIPP’D 28.28 31.20
2hr TSI ASI 24.95 35.81
SKIPP’D 30.38 38.16
ASI TSI 27.42 35.31
SKIPP’D 26.17 35.01
3hr TSI ASI 29.86 38.02
SKIPP’D 31.80 39.12
ASI TSI 30.04 36.61
SKIPP’D 29.61 41.13
4hr TSI ASI 34.37 41.27
SKIPP’D 36.60 41.34
ASI TSI 35.71 38.67
SKIPP’D 36.16 45.28
Table 14. Forecasting Performance with 16 weeks of training.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
1hr TSI ASI 22.76 28.97
SKIPP’D 33.12 36.93
ASI TSI 23.33 35.07
SKIPP’D 25.74 31.01
2hr TSI ASI 25.30 31.55
SKIPP’D 30.75 38.18
ASI TSI 27.48 36.57
SKIPP’D 25.83 32.22
3hr TSI ASI 28.86 36.28
SKIPP’D 33.10 39.83
ASI TSI 31.92 39.46
SKIPP’D 31.04 37.69
4hr TSI ASI 33.99 41.36
SKIPP’D 38.20 42.66
ASI TSI 37.50 42.14
SKIPP’D 38.25 42.46
We conducted a series of experiments to assess the impact of training data size on model performance during fine-tuning. We utilized training splits of 2, 4, 8, 12, and 16 weeks of data at the new site. For each training duration, we performed experiments with different random splits of the corresponding number of weeks and reported the results accordingly.
The results are presented in Tables 10, 11, 12, 13, and 14. Figure 3 was constructed by systematically aggregating the results from our fine-tuning experiments, encapsulating the performance trends observed across different training durations. By leveraging visualization techniques, the figure provides a holistic representation of how the model adapts as more site-specific data becomes available. It effectively summarizes variations in performance across different random splits of training data and across different sets of source and target datasets.
We employed 95% confidence intervals for all experiments, spanning diverse transfer learning settings and random sampling of the fine-tuning data. To rigorously compare our method with the baseline across different weekly intervals, we applied a paired t-test at a significance level of 0.001 (i.e., less than a 0.1% chance of incorrectly rejecting the null hypothesis). In every instance, the observed p-values fell below this threshold, demonstrating that SPIRIT achieves statistically significant performance improvements over the baseline.

### v2

@@
-SPIRIT: Short-term Prediction of solar IRradIance for zero-shot Transfer learning using Foundation Models
-Abstract.
+SPIRIT: Short-Term Prediction of Solar IRradIance for Transfer Learning Using Foundation Models
+Abstract
 Traditional solar forecasting models are based on several years of site-specific historical irradiance data, often spanning five or more years, which are unavailable for newer photovoltaic farms.
-As renewable energy is highly intermittent, building accurate solar irradiance forecasting systems is essential for efficient grid management and enabling the ongoing proliferation of solar energy, which is crucial to achieve the United Nations’ net zero goals.
-In this work, we propose SPIRIT, a novel approach leveraging foundation models for solar irradiance forecasting, making it applicable to newer solar installations.
-Our approach outperforms state-of-the-art models in zero-shot transfer learning by about 70%, enabling effective performance at new locations without relying on any historical data.
+As renewable energy is highly intermittent, building accurate solar irradiance forecasting systems that are data-efficient is essential for efficient grid management and enabling the ongoing proliferation of solar energy, which is crucial to achieve the United Nations’ net zero goals.
+In this work, we propose SPIRIT, a novel framework leveraging foundation models for solar irradiance forecasting, making it applicable to newer solar installations.
+Our approach outperforms state-of-the-art models in zero-shot transfer learning by upto 70%, enabling effective performance at new locations without relying on any past data.
 Further improvements in performance are achieved through fine-tuning, as more location-specific data becomes available.
 These findings are supported by statistical significance, further validating our approach.
-SPIRIT represents a pivotal step towards rapid, scalable, and adaptable solar forecasting solutions, advancing the integration of renewable energy into global power systems.
-Solar Forecasting, Renewable Energy, Foundation Models, Transfer Learning, Zero-shot Learning, Fine-tuning, Deep Learning
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Computing methodologies Machine learningccs: Applied computing Forecastingccs: Hardware Renewable energyccs: Computing methodologies Foundation modelsccs: Computing methodologies Transfer learning
-Figure 1.
-Illustration of our system: A vision encoder (top-left) extracts embeddings from a sky camera image sampled from a diverse set spanning multiple locations and setups.
+By dramatically reducing the forecasting setup timeline, SPIRIT accelerates solar farm deployment in all potential global sites, most of which lack historical data, thereby democratizing access to clean energy and enabling participation in the renewable energy transition.
+Codehttps://github.com/surya-ravindra06/SPIRIT-Official
+Figure 1: Illustration of our system: A vision encoder (top-left) extracts embeddings from a sky camera image sampled from a diverse set spanning multiple locations and setups.
 Physics-inspired features are derived and integrated with auxiliary values, then merged with the image embedding (top-middle) into a unified representation.
 For nowcasting (right), a regressor predicts Global Horizontal Irradiance from this feature vector.
-For forecasting (bottom), a time-series model processes past feature vectors to create a context embedding, which is concatenated with a future covariate vector—constructed from known future values—to form the final latent representation.
+For forecasting (bottom), a time-series model processes past feature vectors to create a context embedding, which is concatenated with a future covariate vector constructed from known future values to form the final latent representation.
 A regressor then maps this representation to future GHI values (bottom-right).
-1. Introduction
-The proliferation of solar energy is paramount for electrification and the global energy transition to meet the Net Zero commitments of the United Nations (Sadhukhan, 2022).
-As the world moves toward renewable sources, solar energy is notable for its accessibility and potential to significantly reduce carbon emissions (Sen, 2008).
-Expanding the solar energy infrastructure is crucial to mitigate the effects of climate change (Bashir et al., 2021) and meet the energy demands arising from sectors such as data centers (Agarwal et al., 2021), transportation (Lee et al., 2016), and buildings (Iyengar et al., 2017).
-Unlike conventional power sources such as thermal and nuclear, solar energy has inherent shortcomings.
-Its intermittency, due to the daily and seasonal variations in sunlight, poses significant challenges for energy grid stability (Abido et al., 2022).
-One notable issue arising from the higher penetration of solar power is the “duck curve” (Iyengar et al., 2016), where the mismatch between solar energy production and peak energy demand leads to significant challenges in grid management.
-Although storage capacity is increasing, electricity grids typically operate as a just-in-time system where energy supply and demand must be balanced (Joskow, 2012).
-To ensure grid efficiency, renewable operators must pay a deviation penalty to discourage unplanned energy contributions, thereby maintaining a balanced and predictable energy supply (Yang et al., 2020).
-Thus, accurate short-term solar predictions are crucial for the efficient operation of the energy grid (Iyengar et al., 2014).
-Existing approaches for short-term forecasting use Sky Cameras — i.e., a fish-eye lens camera positioned to look directly towards the zenith — which require extensive site-specific data to train models (Hammond et al., 2024; Gao and Liu, 2022).
-These models have demonstrated the ability to develop high-accuracy models, albeit using training data spanning multiple years.
-With the overall solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030 (isa, 2023), 90% of the solar farms worldwide will have negligible data to train custom models from scratch.
-Thus, lack of sufficient site-specific solar data underscores the need for approaches that do not compromise model performance.
-With the advent of vision foundation models, we have seen improvement in accuracy of various Computer Vision tasks — such as feature extraction, object detection, etc.
-— using zero-shot and few-shot approaches (i.e., with limited or no custom training data) (Dosovitskiy, 2020; Zohar et al., 2023; Jeeveswaran et al., 2022).
-In addition, physics-inspired feature engineering has significantly improved model performance by incorporating domain-specific knowledge, leading to more accurate and interpretable predictions in real-world problems (Ompusunggu and Hostens, 2021; Erdmann et al., 2020).
-In this work, our hypothesis is as follows: Can we leverage state-of-the-art vision foundation models and physics-inspired features, along with transfer learning strategies, to reduce the dependence on site-specific sky camera imagery data?
-To address these challenges, we introduce SPIRIT, a novel approach to solar irradiance forecasting with an inductive bias toward enhanced generalizability.
-In designing, implementing, and evaluating our approach, we make the following contributions:
-(1) We develop a novel system that leverages foundation models and physics-informed features, eliminating the need for site-specific model training while enabling effective adaptation across diverse transfer learning scenarios.
-The flexibility of our framework ensures seamless integration of future advancements in vision models without requiring significant architectural modifications.
+1 Introduction
+The proliferation of solar energy is paramount for electrification and the global energy transition to meet the Net Zero commitments of the United Nations (sadhukhan2022net).
+Solar energy is notable for its potential to reduce carbon emissions (sen2008solar).
+Expanding the solar energy infrastructure is crucial to mitigate the effects of climate change (bashir2021enabling) and meet the energy demands arising from sectors such as data centers (agarwal2021redesigning), transportation (lee2016shared), and buildings (iyengar2017cloud).
+Unlike conventional sources, solar energy presents unique challenges due to its intermittency from daily and seasonal sunlight variations, significantly impacting grid stability (grid_management).
+High solar penetration creates the ”duck curve” phenomenon (iyengar2016analyzing), where misalignment between solar production and peak demand complicates grid management.
+Despite increasing storage capacity, electricity grids function primarily as just-in-time systems requiring balanced energy supply and demand (joskow2012creating).
+To maintain grid reliability, operators face deviation penalties for unplanned energy contributions (yang2020penalty), necessitating accurate short-term solar predictions for efficient grid operation (iyengar2014solarcast).
+Existing approaches for short-term forecasting use sky cameras, a fish-eye lens positioned to look directly towards the zenith, which require extensive site-specific data to train models (hendrikx2024_short_term_forecast; wacv2022).
+While these methods have demonstrated high accuracy, they rely on multi-year training datasets.
+With global solar photovoltaic (PV) capacity projected to expand from 1 terawatt in 2022 to 10 terawatts by 2030 (isa2023), 90% of the solar farms worldwide will have negligible data to train custom models from scratch.
+Thus, lack of site-specific data underscores the need for novel approaches.
+Vision foundation models have revolutionized numerous computer vision tasks such as feature extraction and object detection, enabling improved accuracy through zero-shot and few-shot approaches with limited data (google_vit; object_detection_foundation_models; vits_gt_cnns1).
+Concurrently, physics-inspired feature engineering has enhanced model performance by integrating domain-specific knowledge, yielding more accurate and interpretable predictions (ompusunggu2021physics; erdmann2020physics).
+Our research hypothesis investigates: Can we leverage state-of-the-art vision foundation models and physics-inspired features, along with transfer learning strategies, to reduce the dependence on site-specific sky camera data?
+To address these challenges, we introduce SPIRIT, a novel solar irradiance forecasting approach with inductive bias toward enhanced generalizability. Our contributions include:
+(1) Development of an innovative system leveraging foundation models, physics-informed features and future covariates that eliminates site-specific training requirements while facilitating adaptation across diverse transfer learning scenarios.
+Our system’s flexibility enables seamless integration of future vision model advancements.
 (2) Motivated by real-world deployment constraints, we demonstrate that SPIRIT can rapidly scale to new solar plant locations without prior sky camera data, significantly accelerating integration into operational workflows.
-2. Related Work
-Traditional methods for solar forecasting have relied heavily on Numerical Weather Prediction (NWP) models and satellite imagery (Markovics and Mayer, 2022).
-While these methods provide valuable insights, they often lack the spatial and temporal resolution required for accurate short-term forecasts.
-For instance, NWP models typically operate on a grid scale of several kilometers and update every few hours which may not capture rapid changes in cloud cover that affect solar irradiance (Kostylev et al., 2011).
-Over the past few years, several time series forecasting approaches have been used for solar forecasting.
-However, they typically operate on a time frame of multiple hours to day-ahead and are not suitable for capturing short-term variations in solar generation due to transient factors such as cloud cover (Iyengar et al., 2014; Falope et al., 2024).
-Use of sky camera imagers for short-term solar forecasting has garnered significant attention in recent years due to their potential to enhance the accuracy of solar power predictions (Hammond et al., 2024; Gao and Liu, 2022; Nie et al., 2024).
-Sky cameras, equipped with fish-eye lenses, capture wide-angle images of the sky, providing valuable data on cloud cover and movement, which are critical factors in solar irradiance forecasting (Dev et al., 2019).
-Recent advancements have focused on leveraging sky cameras to address the limitations of traditional approaches.
-Hammond et al.
-(Hammond et al., 2024) and Gao et al.
-(Gao and Liu, 2022) demonstrated the potential of sky cameras in developing high-accuracy models for short-term solar forecasting.
-These studies utilized extensive site-specific data collected over multiple years to train their models, achieving significant improvements in forecast accuracy compared to traditional methods.
-Siddiqui et al.(Siddiqui et al., 2019) proposed a deep learning framework using sky-camera images and auxiliary meteorological data to predict solar irradiance.
-Their approach employs a convolutional neural network (CNN) with dilated convolutions, followed by an LSTM for temporal forecasting up to four hours ahead.
-By training on 10 years of data, they demonstrated that incorporating auxiliary data such as temperature, wind speed, and relative humidity enhances generalization and stability in predictions.
-Similarly, Gao et al.
-(Gao and Liu, 2022) introduced a transformer-based architecture that integrates a clear sky model to estimate the residual irradiance beyond clear-sky assumptions.
-Trained on 10 years of data, their model achieves improved forecasting accuracy compared to earlier CNN-LSTM-based methods.
-Both works underscore the importance of leveraging sky images and auxiliary data for precise solar nowcasting and forecasting.
-Despite their promise, sky camera-based approaches face challenges related to data availability.
-With the global solar PV fleet expected to increase from 1 TW in 2022 to 10 TW by 2030, a large majority of the solar farms worldwide will have negligible historical data to train custom models from scratch.
-Building upon these challenges, it becomes evident that addressing the limited availability of site-specific data is critical for advancing solar forecasting.
-Although the use of sky cameras and auxiliary data has substantially improved short-term predictions, the scalability of these methods remains constrained by the dearth of historical data at many solar installations.
-In this context, transfer learning emerges as a promising solution, as it enables the leveraging of knowledge from pre-trained models and the adaptation of learned representations across different datasets and locations.
-Notably, previous work such as Nie et al(Nie et al., 2024) has demonstrated that training on a fusion of multiple datasets yields models that perform better on each individual dataset, thereby highlighting the potential benefits of cross-dataset knowledge transfer.
-3. SPIRIT Design
-3.1. Key Concepts and Problem Setup
-Nowcasting refers to the prediction of solar power generation over very short time horizons, typically ranging from a few minutes to a few hours (Lee et al., 2017).
-In contrast, short-term forecasting extends the prediction horizon to cover periods from one hour to 24 hours (Remund and Müller, 2012).
-Methods developed to provide forecasts utilize various data sources, such as satellite data (Lopes et al., 2021; Lee et al., 2017), weather station observations (Lee et al., 2017), and sky camera images (Gao and Liu, 2022; Xu et al., 2015; Siddiqui et al., 2019).
-Nowcasting and short-term forecasting are indispensable for managing the intermittency of solar power, allowing grid operators to perform better scheduling, dispatching, and balancing of energy resources (Dairi et al., 2020; Aouidad and Bouhelal, 2024).
+2 Related Work
+Traditional solar forecasting methods primarily utilize Numerical Weather Prediction models and satellite imagery (markovics2022comparison).
+These approaches often lack sufficient spatial and temporal resolution for accurate short-term forecasts.
+NWP models typically operate at kilometer-scale grids with updates every few hours, failing to capture rapid cloud cover changes affecting solar irradiance (kostylev2011solar).
+Recent years have seen various time series forecasting approaches for solar prediction, yet these generally function at multi-hour to day-ahead timescales, rendering them inadequate for detecting short-term solar generation variations from transient factors like cloud cover (iyengar2014solarcast).
+Sky cameras have gained prominence due to their potential to enhance short-term solar power prediction (wacv2022; yuhao_transfer_learning).
+These fish-eye lenses capture wide-angle sky images that provide crucial data on cloud cover and movement, key determinants of solar irradiance forecasting (dev2019estimating).
+Recent research has focused on utilizing sky cameras to overcome prior limitations.
+wacv2022 and hammond2024 demonstrated sky cameras’ efficacy in developing high-precision short-term solar forecasting models.
+Their studies employed extensive site-specific data collected across multiple years, yielding models with significantly improved forecast accuracy compared to conventional methods.
+talha2019 developed a deep learning framework combining sky-camera imagery with meteorological data for solar irradiance prediction.
+Their approach integrates a CNN and temporal LSTM for forecasts up to four hours ahead, demonstrating that auxiliary data (temperature, wind speed, relative humidity) enhances prediction generalization when trained on 10 years of data.
+wacv2022 proposed a transformer-based architecture incorporating a clear sky model to estimate residual irradiance, achieving superior accuracy compared to CNN-LSTM methods using 10 years of training data.
+Both studies emphasize the efficacy of integrating sky imagery with auxiliary data for precise solar forecasting.
+However, these approaches face data availability constraints, particularly significant as global solar PV capacity expands from 1 TW (2022) to a projected 10 TW by 2030, leaving most solar installations around the world with insufficient historical data for custom model development.
+Limited availability of site-specific data presents a critical challenge for advancing solar forecasting.
+Despite significant improvements in short-term predictions through sky cameras and auxiliary data integration, these methods face scalability constraints due to insufficient historical data at most solar installations.
+Transfer learning emerges as a promising solution, enabling knowledge leveraging from pre-trained models and adaptation of learned representations across diverse datasets and locations.
+Recent research (yuhao_transfer_learning; paletta_transfer_learning) demonstrates that multi-dataset fusion training and data augmentation methods yield models with superior performance, highlighting cross-dataset knowledge transfer benefits in solar forecasting.
+3 Design and Implementation
+3.1 Key Concepts
+Nowcasting and Forecasting: Nowcasting, the prediction of solar power generation over very short time horizons, typically ranges from a few minutes to a few hours (nowcasting_defn).
+In contrast, short-term forecasting extends the prediction horizon to cover periods from one hour to 24 hours (forecasting_defn).
+Methods developed to provide forecasts utilize various data sources, such as satellite data (nowcasting_satellite; nowcasting_defn), weather station observations (nowcasting_defn), and sky camera images (wacv2022; talha2019).
+Nowcasting and short-term forecasting are indispensable for managing the intermittency of solar power, allowing grid operators to perform better scheduling, dispatching, and balancing of energy resources (aouidad2024machine).
 Sky Camera: Sky cameras enhance nowcasting and short-term forecasting by capturing sky images with fish-eye lenses, providing detailed cloud movement, and sun position data.
-These images enable algorithms to track cloud dynamics and predict their trajectories, essential for estimating solar irradiance (Saraswat et al., 2023; Dev et al., 2019).
+These images enable algorithms to track cloud dynamics and predict their trajectories, essential for estimating solar irradiance (saraswat2023sky; dev2019estimating).
 Offering a low-latency alternative to weather satellites, sky cameras facilitate real-time monitoring.
-However, variations in camera setup and quality affect image appearance, as shown in Figure 4 in Appendix A.1.
-As a key tool in solar forecasting, sky cameras contribute to more reliable energy predictions (Rajagukguk et al., 2021).
-Further details are provided in Appendix B.
-Irradiance measurements: Understanding solar irradiance requires distinguishing between three key measurements:
-(1) Direct Normal Irradiance (DNI): The amount of solar radiation received per unit area on a surface perpendicular to the sun’s rays without being scattered or diffused by the atmosphere.
-(2) Diffuse Horizontal Irradiance (DHI): The portion of solar radiation that reaches a horizontal surface after being scattered by molecules, aerosols, and clouds in the atmosphere.
-Unlike DNI, DHI comes from all directions in the sky and plays a crucial role during overcast conditions when direct sunlight is obstructed.
-(3) Global Horizontal Irradiance (GHI): The total solar radiation received on a horizontal surface, combining both direct and diffuse components.
-GHI is the sum of DNI, projected onto a horizontal plane, and DHI:
-(1) $GHI=DNI\times\cos(\theta)+DHI$
-where $\theta$ is the angle between the direction of incoming solar radiation and the vertical, called the zenith angle.
+However, variations in camera setup and quality affect image appearance, as shown in Figure 4 in the Appendix.
+As a key tool in solar forecasting, sky cameras contribute to more reliable energy predictions (rajagukguk2021deep).
+Further details are provided in Appendix C.
+Solar Irradiance necessitates distinguishing between three fundamental measurements.
+Direct Normal Irradiance (DNI) quantifies solar radiation received per unit area perpendicular to the sun’s rays without atmospheric scattering.
+Diffuse Horizontal Irradiance (DHI) measures scattered radiation reaching a horizontal surface from all sky directions, becoming crucial during cloud cover.
+Global Horizontal Irradiance (GHI), the total radiation on a horizontal surface, encompasses both of these components and is mathematically expressed as:
+$GHI=DNI\times\cos(\theta)+DHI$ (1)
+where $\theta$ represents the zenith angle, defined as the angle between the direction of incoming solar radiation and the vertical.
 GHI is the most commonly used irradiance measure in solar energy applications, as it directly influences photovoltaic (PV) panel performance and solar power generation, making it the primary focus of research in irradiance forecasting.
 Henceforth, unless explicitly stated otherwise, any mention of irradiance or solar irradiance refers specifically to Global Horizontal Irradiance.
-Photovoltaic Power Output: PV power output refers to the electricity generated by solar panels from incoming solar radiation.
-While it is primarily driven by GHI (Vilanova et al., 2020), factors like temperature, and system losses also play a role.
-Under stable conditions, the relationship between GHI and PV output is roughly linear (Razak et al., 2016; Natheer Tuaimah and Al-Saidi, 2019).
+Photovoltaic Power Output refers to the electricity generated by solar panels from incoming solar radiation.
+While it is primarily driven by GHI (ghi_pv_linear), factors like temperature, and system losses also play a role.
+Under stable conditions, the relationship between GHI and PV output is roughly linear (pv_output).
 Since PV output is a more actionable metric for grid management and energy planning, predicting it directly is often more desirable.
-3.2. Nowcasting Architecture
-We propose an architecture that encodes sky images into vector representations, which are augmented with auxiliary data and physics-based features.
-This representation captures information about the GHI, which is then effectively extracted by a regression model.
+3.2 Physics-inspired Feature Engineering
+Clear sky models (ineichen1; clearsky1) are mathematical models that estimate the solar irradiance at a given location under cloud-free conditions, serving as a representation of the maximum possible radiation reaching the Earth’s surface.
+These models leverage fundamental atmospheric physics and employ mathematical formulations based on solar geometry, atmospheric transmittance, and radiative transfer (clearsky1).
+The Ineichen model (ineichen1) requires inputs such as latitude, longitude, time, and date, which are readily available.
+This allows clear sky irradiance values to be readily computed and incorporated into our model as features, providing a reference for expected irradiance levels in the absence of cloud interference.
+Physics behind solar irradiance: Solar irradiance varies significantly based on site-specific parameters like panel orientation and solar positioning.
+The critical angle of incidence ($\theta$), representing the angle between incoming solar rays and panel normal, is calculated as $\cos(\theta)=\cos(\theta_{z})\cdot\cos(\beta)+\sin(\theta_{z})\cdot\sin(\beta)\cdot\cos(\gamma-\alpha)$, where $\theta_{z}$ and $\gamma$ denote solar zenith and azimuth angles, while $\beta$ and $\alpha$ represent panel tilt and azimuth angles, respectively.
+These physics-inspired features enable the creation of future covariates through their deterministic calculation capabilities independent of temporal constraints, thereby enhancing time-series information for predictive modeling.
+3.3 Nowcasting Architecture
+We propose an architecture that encodes sky images into vector representations, which are augmented with auxiliary data and physics-based features to capture GHI information that is then effectively extracted by a regression model.
 Let $\mathcal{X}$ be the set of sky camera images, and $\mathcal{D}$ be the dataset, defined as $\mathcal{D}=\{(X_{i},\mathbf{A}_{i},y_{i})\}_{i=1}^{N}$, where $X_{i}\in\mathcal{X}$ represents the $i$-th sky image, $\mathbf{A}_{i}\in\mathbb{R}^{k}$ corresponds to the auxiliary features such as azimuth and zenith angles of the Sun, and $y_{i}\in\mathbb{R}^{+}$ are the corresponding solar irradiance measurements.
@@
 These features incorporate established solar engineering principles, such as clear sky irradiance, and panel tilt and orientation, as defined in Subsection 3.2.
-The feature vector $\mathbf{P}$ is given by:
-$\mathbf{P}\in\mathbb{R}^{p}$
-where $p$ represents the number of physics-based features extracted from the auxiliary data.
+The feature vector is given by $\mathbf{P}\in\mathbb{R}^{p}$ where $p$ represents the number of physics-based features extracted from the auxiliary data.
 The final feature representation $\mathbf{f}\in\mathbb{R}^{d+k+p}$ is constructed by concatenating the image embedding $\mathbf{Z}$, raw auxiliary measurements $\mathbf{A}$, and the physics-based features $\mathbf{P}$:
 $\mathbf{f}=\mathbf{Z}\oplus\mathbf{A}\oplus\mathbf{P}$
-where $\oplus$ denotes the concatenation operation.
-This combined representation leverages data-driven features, visual features, and domain-specific engineering knowledge, providing a comprehensive characterization of each sample $(X_{i},\mathbf{A}_{i},y_{i})\in\mathcal{D}$.
+where $\oplus$ denotes concatenation.
+This combined representation leverages data-driven, visual and domain-specific features, providing a comprehensive characterization of each sample $(X_{i},\mathbf{A}_{i},y_{i})\in\mathcal{D}$.
 A regression function $R_{\omega}:\mathbb{R}^{d+k+p}\rightarrow\mathbb{R}^{+}$, parameterized by weights $\omega$, is defined such that:
 $\hat{y}=R_{\omega}(\mathbf{f})=R_{\omega}(E(X)\oplus\mathbf{A}\oplus\mathbf{P})$
-Nowcasting loss function $\mathcal{L}_{nowcast}(\omega)$ is defined as the average of the individual regression losses for each sample, where each individual loss measures the discrepancy between the predicted $\hat{y}_{i}=R_{\omega}(\mathbf{f}_{i})$ and the true value $y_{i}$:
+Nowcasting loss function $\mathcal{L}_{nowcast}(\omega)$ is defined as the average of the individual regression losses for each sample, where each individual loss measures the discrepancy between the predicted $\hat{y}_{i}=R_{\omega}(\mathbf{f}_{i})$ and the true value $y_{i}$ across all $N$ samples in the dataset:
 $\mathcal{L}_{nowcast}(\omega)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(R_{\omega}(\mathbf{f}_{i}),y_{i})$
@@
 To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{nowcast}(\omega)$ using gradient-based methods.
-3.3. Forecasting Architecture
+3.4 Forecasting Architecture
 Our forecasting architecture processes sequences of sky images to predict GHI across multiple future intervals.
-Each image is encoded using the embedding and augmentation approach from Section 3.2.
-A time-series model captures a latent representation of past features, while predictable future covariates, such as the zenith angle, are precomputed and integrated as a vector.
-The combined past and future representations are then input into a regressor to generate GHI predictions.
+Each image is encoded using the embedding and augmentation approach from Section 3.3, creating a foundation for temporal analysis.
 A sequence of $T$ images $X_{1:T}=\{X_{1},X_{2},\dots,X_{T}\}$ along with their corresponding auxiliary features $\mathbf{A}_{1:T}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{T}\}$, where each $\mathbf{A}_{t}\in\mathbb{R}^{k}$ represents the auxiliary feature vector at time $t$, is given.
@@
 where $\oplus$ denotes concatenation, providing a comprehensive characterization of each sample $(X_{t},\mathbf{A}_{t},y_{t})\in\mathcal{D}$.
-Thus, the collection of feature vectors over the sequence of $T$ time steps is given by:
-$\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$
-where $\mathbf{F_{1:T}}$ represents the set of concatenated feature representations created for each timestamp in the sequence.
-Given the collection of feature vectors $\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$, a time-series model $\mathcal{M}$ is used to encode the observed sequence into a latent vector $\mathbf{L}\in\mathbb{R}^{m}$, which captures the full context of the input data series while retaining its temporal patterns and dependencies:
+Thus, the set of feature vectors for each timestamp over the sequence of $T$ time steps is given by $\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$.
+A time-series model $\mathcal{M}$ is used to encode the observed sequence $\mathbf{F_{1:T}}$ into a latent vector $\mathbf{L}\in\mathbb{R}^{m}$, which captures the full context of the input data series while retaining its temporal patterns and dependencies:
 $\mathbf{L}=\mathcal{M}(\mathbf{F_{1:T}})\in\mathbb{R}^{m}$
-where $\mathcal{M}$ represents the time-series model that transforms the observed sequence of feature vectors $\mathbf{F_{1:T}}$ into a compact representation in the latent space $\mathbb{R}^{m}$.
-To integrate known future information, derived from the spatiotemporal context of time and location, future covariate vectors $\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$ are constructed for each forecast time $T+\tau_{i}$, .
-The full covariate vector $\mathbf{C}\in\mathbb{R}^{q\cdot H}$ is then formed by concatenating these individual representations across all forecast horizons:
-$\mathbf{C}=\bigoplus_{i=1}^{H}\mathbf{C}_{T+\tau_{i}},\quad\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$
-We concatenate the future covariate vector $\mathbf{C}$ with the latent representation of the past time steps $\mathbf{L}$, forming the final vector that encompasses all relevant information:
-$\mathbf{h}=\mathbf{L}\oplus\mathbf{C}$
+To integrate known future information, derived from the spatiotemporal context of time and location, future covariate vectors $\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$ are constructed for each forecast time $T+\tau_{i}$.
+The full covariate vector $\mathbf{C}\in\mathbb{R}^{q\cdot H}$ is then formed by concatenating these individual representations across all $H$ forecast horizons:
+$\mathbf{C}=\mathop{\mathchoice{\displaystyle\oplus}{\textstyle\oplus}{\scriptstyle\oplus}{\scriptscriptstyle\oplus}}_{i=1}^{H}\mathbf{C}_{T+\tau_{i}},\quad\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$
+We concatenate the future covariate vector $\mathbf{C}$ with the latent representation of the past time steps $\mathbf{L}$, forming the final vector $\mathbf{h}$ that encompasses all relevant information: $\mathbf{h}=\mathbf{L}\oplus\mathbf{C}$.
 This ensures that both past contextual information as well as known future data contribute to the forecasting process.
-Next, a regression function $R_{\omega}:\mathbb{R}^{m+q\cdot H}\to\mathbb{R}^{H}$, parameterized by $\omega$, is applied to the vector $\mathbf{h}\in\mathbb{R}^{m+q\cdot H}$ to generate the corresponding predicted GHI values.
+A regression function $R_{\omega}:\mathbb{R}^{m+q\cdot H}\to\mathbb{R}^{H}$, parameterized by $\omega$, is applied to the vector $\mathbf{h}\in\mathbb{R}^{m+q\cdot H}$ to generate the corresponding predicted GHI values.
 The regressor outputs a vector $\hat{\mathbf{y}}\in\mathbb{R}^{H}$ of predicted GHI values for the forecast time intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$:
@@
 where each $\hat{y}_{i}$ corresponds to the irradiance forecast for the time interval $T+\tau_{i}$, and the vector $\hat{\mathbf{y}}$ represents the full set of predicted irradiance values across all forecast intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$.
-Forecasting loss function $\mathcal{L}_{forecast}(\omega)$ is defined as the mean of the individual regression losses computed over all forecast intervals $T+\tau_{j}$ for each sample $i$ Specifically, the total loss is given by:
+Forecasting loss function $\mathcal{L}_{forecast}(\omega)$ is defined as the mean of the individual regression losses computed over all forecast intervals $T+\tau_{j}$ for each sample $i$ across $N$ samples and $H$ forecast horizons.
+The total loss is given by:
 $\mathcal{L}_{forecast}(\omega)=\frac{1}{N\cdot H}\sum_{i=1}^{N}\sum_{j=1}^{H}\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$
-where $\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$ is the individual regression loss for the forecast interval $T+\tau_{j}$ for sample $i$ To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{forecast}(\omega)$ using gradient-based optimization methods.
+where $\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$ is the individual regression loss for the forecast interval $T+\tau_{j}$ for sample $i$ .
+To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{forecast}(\omega)$ using gradient-based optimization methods.
 The complete architecture is illustrated in Figure 1.
-The Significance of Generalized Encoders: A key distinction of our approach is that in prior work (Gao and Liu, 2022; Hasan, 2023; Siddiqui et al., 2019), the encoder $E$ is a vision model typically trained on data from a specific location and camera setup.
-Furthermore, studies aiming for generalizability typically rely on training models using a fusion of solar datasets from multiple locations (Nie et al., 2024; Despotovic et al., 2024).
+The Significance of Generalized Encoders: A key distinction of our approach is that in prior work (wacv2022; prior_methods_site_specific; talha2019), the encoder $E$ is a vision model typically trained on data from a specific location and camera setup.
+Furthermore, studies aiming for generalizability typically rely on training models using a fusion of solar datasets from multiple locations (yuhao_transfer_learning; site_specific_trained_transfer_learning).
 In contrast, we argue, and later demonstrate, that leveraging a foundation model, a highly generalizable feature extractor, provides a more robust $E$ function.
 A foundation model not only matches the performance of site-specific encoders at a given location with a particular setup but also demonstrates an unparalleled advantage in generalizing across diverse locations and camera setups.
-3.4. Physics-inspired Feature Engineering
-Clear sky models (Ineichen and Perez, 2002; Stein et al., 2012; Perez et al., 2002; Mueller et al., 2004) are mathematical models that estimate the theoretical solar irradiance at a given location under cloud-free conditions, serving as a representation of the maximum possible radiation reaching the Earth’s surface.
-These models leverage fundamental atmospheric physics and employ mathematical formulations based on solar geometry (Stein et al., 2012), atmospheric transmittance (Stein et al., 2012), and radiative transfer (Stein et al., 2012) to derive estimations of GHI, DNI and DHI under clear sky conditions.
-The Ineichen clear sky model (Ineichen and Perez, 2002) requires inputs such as latitude, longitude, time, and date, which are readily available.
-This allows clear sky irradiance values to be readily computed and incorporated into our model as features, providing a reference for expected irradiance levels in the absence of cloud interference.
-Physics behind solar irradiance: Solar irradiance, the power per unit area received from the Sun in the form of electromagnetic radiation, is measured in watts per square meter $(W/m^{2})$.
-The amount of solar irradiance received by a solar panel depends on additional site-specific factors, including the panel’s tilt and orientation angle, the Sun’s altitude and azimuth, and the geographic location’s latitude and longitude.
-We first look at the angle of incidence ($\theta$(Laboratories, [n.
-d.]), i.e.
-the angle between the incoming solar rays and the normal to the surface of the solar panel.
-It can be calculated using the following formula:
-(2) $cos(\theta)=cos(\theta_{z})\cdot cos(\beta)+sin(\theta_{z})\cdot sin(\beta)\cdot cos(\gamma-\alpha)$
-where $\theta_{z}$ and $\gamma$ are the solar zenith and azimuth angles respectively. While $\beta$ and $\alpha$ are the tilt and azimuth angles of the panel.
-We calculate the effective irradiance by adding the three main components: direct, diffuse, and reflected irradiance (see below):
-(3) $I_{panel}=DNI\cdot cos(\theta)+DHI\cdot\frac{1+cos(\beta)}{2}+GHI\cdot\rho\cdot\frac{1-cos(\beta)}{2}$
-where $I_{panel}$ is the effective irradiance and $\rho$ is the ground reflectance.
-4. SPIRIT Implementation
-4.1. Nowcasting
-In our approach, we utilize the pre-trained Google Vision Transformer (ViT) (Dosovitskiy, 2020), a model with 632 million parameters, to generate embeddings for sky camera images.
-To reduce sensor dependence and focus on image features, we exclude meteorological sensor data, incorporating only auxiliary variables such as zenith and azimuth angles, clear sky irradiance, panel tilt, and orientation.
-These image embeddings are subsequently concatenated with the auxiliary vector to form the final feature representation.
-The combined feature vectors, paired with their corresponding ground truth GHI values, are then used to train an XGBoost regressor within a supervised learning framework.
-The model is optimized by minimizing the Mean Squared Error (MSE) loss function, which measures the difference between the predicted and actual GHI values.
-4.2. Forecasting
-For forecasting, we employ the Google Vision Transformer (ViT) (Dosovitskiy, 2020) to generate image embeddings, which are subsequently concatenated with the auxiliary variables to form a comprehensive feature representation.
-To account for temporal dependencies, we input a sequence of six images, representing a 1-hour context window, into a transformer-based time-series encoder (Vaswani, 2017).
-This encoder processes the temporal sequence and learns a latent representation of the past context, which is then fused with a future covariate vector that includes azimuth and zenith angles, as well as clear sky GHI.
-The resulting representation is passed through a multi-layer perceptron (MLP) to predict the solar irradiance for the 1-hour, 2-hour, 3-hour, and 4-hour forecast intervals.
-This implementation exemplifies one approach in our framework, with additional variations incorporating different vision encoders of varied sizes in the ablation studies detailed in Section 7.
-5. Evaluation Methodology
-5.1. Datasets
-We evaluate our methods using three publicly available datasets: TSI880 (Andreas and Stoffel, 1981), ASI16 (Andreas and Stoffel, 1981), and SKIPP’D (Nie et al., 2023).
+3.5 Implementation
+Our implementation leverages the pre-trained Google ViT (google_vit) for image embedding generation, deliberately excluding all meteorological sensor data to significantly reduce sensor dependence.
+Instead, we incorporate only the zenith and azimuth angles and, physics-based features (clear sky irradiance values, panel tilt and orientation).
+These carefully selected image embeddings are concatenated with the auxiliary vector to form robust feature representations that train an XGBoost regressor for nowcasting.
+For forecasting, we extend these feature representations across a one-hour context window (six consecutive timestamps) and process them through a transformer encoder (vaswani2017attention).
+The resulting temporal representation is fused with future covariates from the auxiliary and physics features before passing through MLP layers to predict solar irradiance at 1-4 hour horizons, using clear-sky GHI as residuals before the final regressor.
+This represents one concrete realization of our system, with comprehensive ablations in Appendix A for understanding the contributions of each component.
+More details on hyperparameters and deployment efficiency are provided in Appendix F.
+4 Evaluation Methodology
+4.1 Datasets and Performance Metrics
+We evaluate our methods using three publicly available datasets: TSI880 (tsi_dataset), ASI16 (tsi_dataset), and SKIPP’D (SKIPPD_dataset).
 The TSI880 and ASI16 datasets, both collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado, provide sky images captured every 10 minutes along with corresponding GHI values and auxiliary data such as air temperature and relative humidity and only differ in camera setup and sensors, with the ASI16 dataset capturing higher-resolution images.
 The SKIPP’D dataset, collected from Stanford University, consists of raw sky images captured every minute and PV power output data, prioritizing finer temporal granularity at the expense of image quality.
-For more details, refer to Appendix A.
+For more details, refer to Appendix B.
 We utilize the TSI880 and ASI16 datasets to investigate the impact of camera setup at the same location.
 To explore location and task shifts, we use the SKIPP’D dataset to evaluate the performance of models trained on GHI data in predicting PV power output.
-The SKIPP’D dataset features lower-resolution images and lacks meteorological data, thereby presenting a more challenging task by limiting the contextual information typically leveraged by prior models (Gao and Liu, 2022; Siddiqui et al., 2019).
+The SKIPP’D dataset features lower-resolution images and lacks meteorological data, thereby presenting a more challenging task by limiting the contextual information typically leveraged by prior models (wacv2022; talha2019).
 To ensure the models learn from higher-quality, information-rich datasets, we train exclusively on the TSI and ASI datasets while evaluation is done across all the datasets, including the more challenging SKIPP’D, allowing us to assess how well the models generalize to lower-quality data and increased domain shifts.
-5.2. Performance Metrics
-We assess the effectiveness of the predicted values using the normalized Mean Absolute Percentage error (nMAP), defined as:
-(4) $\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$
+We assess the effectiveness of the predicted values using the normalized Mean Absolute Percentage error (nMAP):
+$\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$ (2)
 where $y_{i}$ represents the actual value and $\hat{y}_{i}$ represents the predicted value for the $i$-th sample, with $i\in\{1,\dots,N\}$.
-It is commonly used for solar irradiance prediction as the normalization ensures that models can be assessed uniformly on datasets with varied value ranges, avoiding biased assessments due to scale differences.
-5.3. Baselines
-To benchmark our proposed method, we compare its performance against the state-of-the-art baseline, Gao et al.
-(Gao and Liu, 2022), who demonstrate state-of-the-art performance for nowcasting and forecasting by training a vision transformer (Dosovitskiy, 2020) from the ground up using 10 years of site-specific data (Andreas and Stoffel, 1981; Gao and Liu, 2022; Siddiqui et al., 2019).
-Their approach for forecasting utilizes a temporal transformer (Vaswani, 2017), also trained on the same duration of data.
-To ensure a fair comparison, we reproduced their architecture and conducted experiments under the same conditions for both Gao et al.’s (Gao and Liu, 2022) model and SPIRIT.
-Table 1.
-Nowcasting performance across multiple datasets: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset for a year are evaluated with nMAP both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017.
-We observe comparable performance when tested in the training setup, but our model demonstrates significantly better zero-shot performance in a new location.
-Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+It is commonly used for solar irradiance prediction as the normalization ensures that models can be assessed uniformly on different datasets, avoiding biased assessments due to scale differences.
+Comprehensive analysis on metrics is in Appendix I.
+4.2 Baselines
+We benchmark against wacv2022, the current state-of-the-art, who achieve superior performance by training a vision transformer (google_vit) and temporal transformer (vaswani2017attention) on 10 years of site-specific data (tsi_dataset; wacv2022; talha2019) with sensor-dependent auxiliary values.
+For equitable evaluation, we reproduced their architecture and conducted experiments under identical conditions for both their model and SPIRIT.
+More details on baselines are in Appendix G.
+Train Test SPIRIT wacv2022
 TSI ASI 27.17 (-62.49) 89.66
@@
 ASI 9.08 (+0.95) 8.13
-Table 2.
-Forecasting performance across multiple datasets and forecast intervals: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset are evaluated with nMAP error both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017 across four forecast intervals: 1hr, 2hr, 3hr, and 4hr.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 1: Nowcasting performance across multiple datasets: SPIRIT and wacv2022 model trained on one dataset for a year are evaluated with nMAP error (lower is better) both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017.
+Performance is comparable when tested in the training setup, but our model demonstrates significantly better zero-shot performance when in a new setup or a new location.
+Interval Train Test SPIRIT wacv2022
 1hr TSI ASI 29.99 (-5.75) 35.74
@@
 ASI 29.29 (-1.73) 31.02
-5.4. Zero-shot Transfer Learning
+Table 2: Forecasting performance across multiple datasets and forecast intervals: SPIRIT and wacv2022 model trained on one dataset are evaluated with nMAP error both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017 across four forecast intervals: 1hr, 2hr, 3hr, and 4hr.
+4.3 Zero-shot Transfer Learning
 To evaluate the zero-shot generalization performance of our models, we analyze two distinct transfer learning scenarios.
-The first scenario examines intra-location generalization, where the models are trained and tested in the same geographic location but under varying camera setups.
-While the environmental conditions remain consistent, variations in camera setup, viewing angles, and image resolutions exist between the training and testing phases.
-When image-based models are trained on data from a particular camera setup, they learn to associate specific regions of the image with key features—such as the position of the sun, cloud formations, or atmospheric conditions—that influence the predicted output.
-However, when the camera setup is altered, the spatial mapping of these features within the image shifts.
-To assess how well the models handle such variations, we train them using the TSI dataset and evaluate them on the ASI dataset, and vice versa.
-The second scenario focuses on cross-location and cross-task generalization, where models trained in one geographic location are tested in another with different environmental and sensor characteristics.
-We train on the TSI and ASI datasets and evaluate on the SKIPP’D dataset, with the task shifting from predicting GHI to PV power output.
-Since GHI and PV output have a nearly linear correlation (Vilanova et al., 2020), this serves as a valid example of heterogeneous transfer learning.
-To account for the significant scale difference between GHI and PV output, model outputs are normalized for comparability.
-We conduct experiments for both nowcasting and forecasting tasks, training the models for one year and testing them on another year to account for seasonal variations, thus ensuring a fair evaluation.
-The nMAP errors are reported in Table 1 for nowcasting and Table 2 for forecasting, comparing SPIRIT with the state-of-the-art in both-the zero-shot transfer learning setups and the traditional settings, where the models are trained and tested using data from the same location and setup, but different years.
-Figure 2.
-We compare the nowcasting performance of SPIRIT and Gao et al.
-using nMAP error.
-The solid lines represent the average performance across different fine-tuning training sizes, measured in weeks of data.
-The shaded regions indicate the xconfidence interval, reflecting variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting fine-tuning data from different starting points throughout the year.
-Figure 3.
-We compare the forecasting performance of SPIRIT and Gao et al.
-using nMAP error across different forecast intervals.
+First, we examine intra-location generalization where models are trained and tested at identical geographic locations but with differing camera configurations.
+Though environmental conditions remain consistent, variations in camera setup alter the spatial mapping of critical features within images.
+Models trained on specific camera setups learn to associate image regions with key elements (sun position, cloud formations, atmospheric conditions) that influence predictions.
+When these spatial relationships shift due to altered camera configurations, model performance deteriorates.
+We evaluate this scenario by training on TSI and testing on ASI (both from Golden, Colorado), and vice versa.
+Second, we investigate cross-location and cross-task generalization, training models in one geographic location and testing in another with different environmental and sensor characteristics.
+Specifically, we train on TSI and ASI datasets and evaluate on SKIPP’D (from Stanford University, California), shifting from GHI prediction to PV power output forecasting.
+Given the nearly linear correlation between GHI and PV output (ghi_pv_linear), this constitutes a valid heterogeneous transfer learning scenario.
+Model outputs are normalized to address scale differences between GHI and PV measurements, as detailed in Appendix F.2.
+For comprehensive evaluation, we conduct experiments for both nowcasting and forecasting tasks, training on one year and testing on another to account for seasonal variations.
+Tables 1 and 2 report nMAP errors for nowcasting and forecasting respectively, comparing SPIRIT against the state-of-the-art in both zero-shot transfer learning and traditional setting (same location and setup).
+Figure 2: Mean nMAP performance of SPIRIT and wacv2022 across varying fine-tuning data sizes in weeks.
+Solid lines denote average accuracy and shaded regions show 95% confidence intervals across multiple runs, including cross-setup (TSI to ASI and vice versa) and cross-location (TSI or ASI to SKIPP’D) transfer scenarios.
+Figure 3: Forecasting performance of SPIRIT and wacv2022 using nMAP error across different forecast intervals.
 Subfigures (a), (b), (c), and (d) correspond to 1-hour, 2-hour, 3-hour, and 4-hour forecasting, respectively.
-The solid lines represent the average performance for each forecast interval, with varying fine-tuning training sizes measured in weeks of data.
-The shaded regions denote the 95% confidence interval, illustrating the variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting weeks of contiguous fine-tuning data from different starting points throughout the year.
-SPIRIT exhibits consistently low variance compared to the baseline, particularly in settings with severely limited data, demonstrating its ability to maintain stability.
-In contrast, the baseline shows high variance, indicating uncertainty in its predictions.
-5.5. Fine-tuning with Limited Data
-Building upon our zero-shot transfer learning experiments, we now investigate the adaptability of our models in a fine-tuning framework, where a limited amount of labeled data from the target domain is available for fine-tuning.
-This scenario closely resembles practical deployment conditions, where prolonged data collection is often infeasible, and models must quickly adapt to new locations with minimal supervision.
-We evaluate transfer learning with limited data in two scenarios: intra-location adaptation and cross-location adaptation, as in Subsection 5.4.
-For both experimental setups, we perform fine-tuning using progressively increasing amounts of labeled data from the target domain—specifically, one, two, three, and four weeks of data from a full year for nowcasting, and two, four, eight, twelve, and sixteen weeks of data for forecasting with testing done on the remaining data from the year.
-Given the greater complexity of forecasting, we extend the fine-tuning experiment to a larger time frame.
-Additionally, due to the requirement for temporal consistency in the time series data, as discussed in Subsection A.2, the number of nowcasting samples for a given time period exceeds that of forecasting samples.
-We implement a selective fine-tuning approach, where only the regressors (see Figure 1) are updated, while the rest of the model is frozen.
-This ensures that the pre-trained feature representations, which capture generalizable spatiotemporal patterns, remain intact while allowing the model to adapt to location- and camera-specific variations.
-As demonstrated in prior work (Nie et al., 2024; Zhou et al., 2020; Sarmas et al., 2022), fine-tuning only the final layers achieves competitive adaptation performance while mitigating the risk of overfitting to the limited target data The nowcasting metrics are shown in Figure 2, and the forecasting metrics are depicted in Figure 3.
-6. Results
-6.1. Zero-shot Transfer Learning
-Tables 1 and 2 present the results for zero-shot transfer learning, demonstrating that our model consistently outperforms the state-of-the-art baseline across both cross-location and cross-setup scenarios in both nowcasting and forecasting tasks.
-When transitioning between camera setups within the same location, our model consistently shows better performance relative to the baseline.
-However, more notably, when moving across different locations, our model achieves up to 45% improvement.
-In this more complex cross-location setting, our model significantly outperforms the baseline, highlighting its superior generalizability and robustness.
-Furthermore, even in the traditional setup where models are trained and tested on the same location, our approach demonstrates enhanced forecasting performance, further emphasizing its effectiveness across diverse deployment conditions.
-6.2. Fine-tuning with Limited Data
-For the analysis of fine-tuning results, we merge cross-setup and cross-location scenarios to ensure a sufficient number of data points for robust confidence interval plots, as depicted in both the nowcasting (Figure 2) and forecasting (Figure 3) tasks.
-Since nowcasting is a relatively simpler task, both models exhibit rapid improvement within the first week.
-However, the baseline model reaches performance saturation early, at approximately 45%, while our model continues to reduce its error, achieving a significant improvement, dropping below 20% within four weeks.
-In forecasting, SPIRIT consistently outperforms the baseline, demonstrating notably lower variance, particularly in data-limited settings (0-2 weeks of data).
+The solid lines represent the average performance, with varying fine-tuning data sizes (in weeks).
+The shaded regions denote the 95% confidence interval, illustrating the variability across multiple runs, including training on one dataset and finetuning and testing on another, as well as utilizing randomized seasonal sampling.
+SPIRIT exhibits consistently better performance and low variance compared to the baseline, particularly with severely limited data, demonstrating its ability to maintain stability over the baseline’s.
+4.4 Fine-tuning with Limited Data
+Following our zero-shot transfer learning experiments, we examine SPIRIT’s adaptability through fine-tuning with limited target domain data.
+This scenario replicates real-world deployment constraints where extensive data collection is impractical and models must adapt rapidly with minimal supervision.
+We evaluate both intra-location and cross-location/cross-task adaptation scenarios as described in Subsection 4.3.
+For both experimental configurations, we incrementally increase the amount of target domain data used for fine-tuning, one to four weeks for nowcasting and two to sixteen weeks for forecasting, with testing performance on the remaining annual data.
+The longer timeframe for forecasting accounts for the reduced sample density due to temporal consistency constraints, highlighted in Appendix B.3, and forecasting’s fundamentally higher complexity relative to nowcasting.
+We employ a selective fine-tuning strategy wherein only the regressor layers (Figure 1) are updated while freezing the remainder of the temporal encoder.
+This preserves the pre-trained feature extractors that capture generalizable spatiotemporal patterns, an essential part of our core hypothesis.
+Selective fine-tuning of terminal layers not only accelerates adaptation but also mitigates overfitting on limited target data (yuhao_transfer_learning; uses_freezing_for_tl; transfer_learning_strategies_comparison).
+Detailed rationale for this choice is provided in Appendix D.3.
+The nowcasting and forecasting performance metrics are presented in Figures 2 and 3, respectively.
+5 Results
+5.1 Zero-shot Transfer Learning
+Tables 1 and 2 present the results for zero-shot transfer learning, demonstrating that our model consistently outperforms the state-of-the-art baseline across both cross-location and cross-setup scenarios in nowcasting and forecasting tasks.
+In nowcasting, SPIRIT exhibits remarkable adaptability when transitioning between camera setups within the same location (TSI to ASI and vice versa) or across different geographical contexts (TSI and ASI to SKIPP’D), outperforming the baseline by up to 70%.
+Similarly, in forecasting tasks, our approach achieves up to 45% improvement when transferring across locations (from TSI/ASI to SKIPP’D).
+These substantial performance gains, particularly in challenging cross-location and cross-task settings, underscore our model’s exceptional generalizability and robustness.
+Furthermore, even in the traditional setup where models are trained and tested on the same dataset and setup, SPIRIT demonstrates enhanced or comparable forecasting performance, further emphasizing its effectiveness across diverse deployment conditions.
+Additional results for our system with different foundation model variants are provided in Appendix E, confirming that the observed improvements stem from SPIRIT’s design.
+5.2 Fine-tuning with Limited Data
+Our experimental analysis merges cross-setup evaluations (training on ASI, finetuning and subsequently testing on TSI, and vice versa) and cross-location scenarios (shifting from ASI or TSI to SKIPP’D) to ensure a sufficient number of experimental instances for the robust confidence interval plots in Figures 2 and 3 for nowcasting and forecasting respectively.
+In the simpler nowcasting task, both models exhibit rapid improvement within the first week.
+However, the baseline reaches performance saturation around 45%, while SPIRIT continues to reduce its error, achieving a significant improvement, dropping below 20% within four weeks.
+In forecasting, SPIRIT consistently outperforms the baseline, demonstrating notably lower variance, particularly in data-limited settings in the first 2 weeks.
 This underscores SPIRIT’s superior stability and reliability, with its nMAP error remaining consistently below that of the baseline.
-In contrast, the baseline model exhibits higher variance, indicating greater inconsistency and confusion in its performance when limited data is available.
+In contrast, the baseline exhibits higher variance, indicating greater inconsistency and confusion in its performance when limited data is available.
 Both models experience a typical performance decline as the forecasting horizon extends from 1-hour to 4-hour forecasts, driven by the increased uncertainty over longer time horizons.
@@
 The transition from a zero-shot configuration to fine-tuning results in noticeable performance improvements; however, the gains diminish after approximately eight weeks of fine-tuning, suggesting that extended fine-tuning beyond this period yields only marginal additional benefits.
-All the results are in Appendix C.
-7. Ablation Studies
-7.1. Investigating Different Vision Encoders
-We examine the impact of different vision models on SPIRIT’s performance, also highlighting the versatility of our system across different foundation models.
-We evaluate the CNN-based ResNet-152 (He et al., 2016), the vision transformer-based DINOv2 Giant (Oquab et al., 2023), and our implementation using Google ViT-Huge (Dosovitskiy, 2020).
-Results summarized in Table 3, demonstrate that the ViT-based models consistently outperform the ResNet-152 CNN model, which can be attributed to the superior capability of ViT architectures in capturing global image features (Jeeveswaran et al., 2022).
-Table 3.
-We explore the impact of using different vision encoders on the overall model performance for nowcasting and forecasting, with training on TSI 2020 and testing on TSI 2021, measured by nMAP error.
-Model Nowcast Forecast
+Detailed results are in Appendix D.
+6 Conclusion
+This work addresses a critical challenge in solar irradiance forecasting: generalizing models to novel geographic locations without historical data.
+SPIRIT leverages pre-trained models and physics-informed features for future covariates, enabling robust performance in data-scarce environments while facilitating post-deployment fine-tuning as site-specific observations accumulate.
+SPIRIT’s modular architecture accommodates emerging vision models, ensuring its methodological relevance while its scalability accelerates the deployment of solar farms in remote and emerging markets, while also enhancing the reliability, cost-effectiveness, and accessibility of solar energy generation to facilitate the global transition toward renewable energy systems.
+7 Future Work and Limitations
+A key limitation is the geographical restriction of datasets to North America due to limited availability of publicly accessible data from other regions, while Southern Hemisphere solar dynamics require further investigation.
+Future work will expand dataset diversity to enhance generalizability.
+Additionally, foundation models introduce computational overhead and inference latency.
+Subsequent research will focus on efficiency optimization for deployment on resource-constrained edge devices.
+Acknowledgements
+We would like to thank the PreCog Lab at IIIT Hyderabad for their valuable guidance and discussions throughout this work.
+In particular, we thank Sreeram Vennam, Jain Hemang Ashok, Vaishnavi Shivkumar, Akshit Sinha, Anish Joishy and Vamshi Krishna Bonagiri for their detailed feedback and thoughtful suggestions during multiple stages of the project.
+We also thank Harpreet Singh, Quentin Paletta, Talha Ahmad Siddiqui, and Tejas Cavale for their valuable inputs and constructive perspectives.
+Finally, we thank the anonymous reviewers for their insightful feedback.
+The views expressed are those of the authors and do not necessarily reflect the official policies or positions of the supporting organizations.
+Appendix A Ablation Studies
+A.1 Impact of the Future Covariates
+We investigate the contribution of the future covariates on SPIRIT’s predictive capacity, by removing them and constraining the model to rely exclusively on historical feature vectors.
+As demonstrated in Table 3, this modification induces performance degradation across all forecasting horizons.
+Error increases of 0.62, 1.01, 1.46, and 2.21 percentage points as the prediction window extends from 1-hour to 4-hour forecasts reveal a clear monotonic trend.
+This progressive deterioration in performance indicates that future covariates provide increasingly valuable information as the prediction horizon extends.
+The pronounced impact on longer-range forecasts demonstrates that future covariates become particularly crucial when predicting under heightened uncertainty, where historical patterns alone prove insufficient for accurate extrapolation.
+Incorporation of future covariates enables the architecture to establish more sophisticated temporal dependencies by providing complementary contextual information that augments the historical patterns encoded in past latent representations, thus enhancing the model’s ability to capture complex dynamics in the underlying time series.
+More details are in Section H.
+Model Configuration Forecast
 +1hr +2hr +3hr +4hr
-ResNet-152 10.50 24.56 27.82 31.23 35.85
-DINOv2 Giant 9.74 21.22 23.56 27.93 33.13
-Google ViT-Huge 9.32 19.96 22.64 26.30 31.58
-7.2. Foundation Model Size
+With Future Covariates 18.96 21.77 25.46 29.89
+Without Future Covariates 19.58 22.78 26.92 32.10
+Table 3: Ablation study comparing the forecasting performance of SPIRIT with and without future covariates across different time horizons.
+We train on TSI 2020 and evaluate on TSI 2021 using nMAP error, where lower values indicate better performance.
+A.2 Investigating Different Vision Encoders
+We evaluate SPIRIT’s performance with diverse vision architectures, including CNN-based (ResNet, EfficientNet, ConvNextv2 (convnext)) and attention-based approaches (DINOv2, Google ViT (google_vit), Apple AIMv2 (aimv2)).
+Results in Table 4 show that transformer architectures yield superior performance compared to traditional convolutional approaches, corroborating findings that self-attention mechanisms better capture the global image features (vits_gt_cnns1) critical for solar forecasting.
+Notably, ConvNeXtv2 matches ViTs in performance, owing to its novel FCMAE framework and GRN mechanism (convnext).
+The architectural modularity of SPIRIT facilitates seamless integration of such emerging models, conferring substantial advantages for sustained operational deployment.
+Model Params Nowcast Forecast
++1hr +2hr +3hr +4hr
+AIMv2 2.72B 9.04 20.89 23.90 27.22 32.71
+DINOv2 1.14B 9.74 21.22 23.56 27.93 33.13
+Google 632M 9.04 18.96 21.77 25.46 29.89
+ConvNeXtv2 88.7M 9.92 19.73 22.09 25.17 29.56
+EfficientNet-B7 66M 13.34 25.20 27.25 29.58 34.27
+ResNet-152 60.3M 10.50 24.56 27.82 31.23 35.85
+Table 4: We use the largest variants of diverse CNN and transformer-based vision encoders to explore their influence on SPIRIT’s nowcasting and forecasting performance, training on TSI 2020 and testing on TSI 2021, measured by nMAP error.
+A.3 Foundation Model Size
 Table 5 presents an analysis of how the size of the foundation model influences the performance of our nowcasting and forecasting architectures.
-Although increasing model size has traditionally been linked to performance gains, we observe that beyond a certain threshold, further scaling yields diminishing returns.
-This suggests that larger models do not always lead to better performance.
-In fact, models with 304M and 86M parameters outperform their larger counterparts with 632M parameters in forecasting and nowcasting, respectively.
-This aligns with recent work, which highlights that adjusting model size based on a computational budget, rather than blindly increasing model size, can lead to more efficient architectures with reduced inference costs (Alabdulmohsin et al., 2023).
-Table 4.
-We evaluate the impact of varying size of the Google ViT vision encoder on the overall performance of the model for both nowcasting and forecasting tasks, with training on TSI 2020 and testing performed on TSI 2021.
-Model Parameters Nowcast Forecast
+Although increasing model size has traditionally been linked to performance gains, we observe that beyond a threshold, further scaling yields diminishing returns or even performance degradation as with DINOv2.
+For transformer-based architectures, models in the moderate parameter range outperform the largest variants, while in the ResNet family, the smallest variant demonstrates superior forecasting performance, whereas the second smallest attains peak nowcasting results.
+These patterns suggest that the impact of model scaling aligns more closely with parameter count than with architectural family distinctions.
+Specifically, for transformer-based models such as DINO and Google-ViT, performance improves consistently with increasing size up to approximately 600M parameters, beyond which gains plateau or regress.
+In contrast, ResNets exhibit distinct scaling dynamics that diverge from those of transformer architectures.
+Model Params Nowcast Forecast
 +1hr +2hr +3hr +4hr
-86M 9.14 21.92 24.07 28.73 34.50
+Google ViT 86M 9.14 21.92 24.07 28.73 34.50
 304M 9.45 19.58 21.95 25.54 30.60
-632M 9.32 19.96 22.64 26.30 31.58
-8. Conclusion
-This work addresses a critical challenge in solar irradiance forecasting: adapting models to new geographic locations with no prior data.
-By utilizing transfer learning and pre-trained models, SPIRIT generalizes well to new locations, reducing the reliance on large, location-specific datasets.
-As more site-specific data becomes available post-deployment, the system can be effectively fine-tuned, improving prediction accuracy and supporting better energy yield estimates and operational planning.
-Additionally, SPIRIT’s modular design allows for the seamless integration of any emerging vision models, ensuring that the framework remains up-to-date with the latest advancements.
-This scalable solution for solar irradiance forecasting can accelerate the deployment of solar farms—particularly in remote and emerging markets.
-SPIRIT supports the transition to renewable energy by enhancing the reliability, cost-effectiveness, and accessibility of solar energy generation.
-9. Future Work and Limitations
-One key limitation is that the datasets used for evaluation are all from North America, largely due to the limited availability of publicly accessible datasets from other regions.
-Specifically, the solar movement patterns and dynamics change in the Southern Hemisphere and need to be studied.
-To improve the generalizability of our system, future work will incorporate data from other continents.
-Additionally, while our model performs well, the use of foundation models introduces real-time inference costs and computational overheads.
-Future efforts will focus on reducing computational efficiency, enabling deployment on resource-constrained edge devices without sacrificing accuracy.
-Appendix A Dataset Details
-Table 5.
-A Comparative Overview of the TSI880, ASI16, and SKIPP’D Datasets: Key Attributes Including Geographical Location, Data Provided, Image Resolution, Collection Frequency, and Annual Sample Size
+632M 9.04 18.96 21.77 25.46 29.89
+DINOv2 22.1M 10.13 20.08 22.53 25.69 29.86
+86.6M 9.56 19.52 21.95 25.00 29.57
+304M 10.04 19.32 21.74 25.29 30.04
+1.14B 9.74 21.22 23.56 27.93 33.13
+ResNet 11.7M 10.47 22.31 24.28 27.28 32.52
+25.6M 10.29 24.36 25.93 29.07 34.10
+44.5M 10.97 24.99 26.21 30.62 34.86
+60.3M 10.50 24.56 27.82 31.23 35.85
+Table 5: Impact of scaling across Google-ViT, DINOv2, and ResNet model families on nowcasting and forecasting, with training on TSI 2020 and testing on TSI 2021, and nMAP as the error metric.
+Results are highlighted for the best performer in each family.
+A.4 Contribution of Auxiliary and Physics Features
+To quantify the individual and combined contributions of different feature modalities, we conduct a systematic ablation study examining the impact of image embeddings, physics-based features, and auxiliary spatiotemporal variables on nowcasting performance.
+As presented in Table 6, the results reveal several critical insights regarding feature importance and complementarity.
+The baseline configuration using only image embeddings achieves a test nMAP of 12.45%, establishing the fundamental predictive capability of visual information.
+The incorporation of either physics features (9.48% nMAP) or auxiliary features (9.55% nMAP) individually provides substantial performance improvements, with both modalities contributing nearly equivalent error reductions.
+Crucially, the combination of physics and auxiliary features without image embeddings yields severely degraded performance (34.23% nMAP), conclusively demonstrating that visual information serves as the primary predictive signal, while physics-based and auxiliary features provide essential complementary enhancements rather than standalone predictive power.
+The optimal configuration integrating all three feature types achieves the lowest error at 9.04% nMAP, validating our feature choice and the importance of our multimodal approach in capturing the complex dynamics inherent in this task.
+Setup nMAP (%)
+Only image embeddings 12.45
+Embeddings + Physics 9.48
+Embeddings + Auxiliary 9.55
+Physics + Auxiliary 34.23
+Embeddings + Physics + Auxiliary 9.04
+Table 6: Ablation on the contribution of image embeddings, physics features, and auxiliary features for nowcasting.
+Results are reported as test nMAP (%, lower is better).
+The model is trained on TSI-2020 and tested on TSI-2021.
+A.5 Variation of Regressors for Nowcasting
+To establish the optimal modeling approach for nowcasting, we conducted a comprehensive evaluation of different regression models in our system while maintaining consistent input features across all models.
+As demonstrated in Table 7, XGBoost achieves superior performance with a test nMAP of 9.04%, substantially outperforming alternative approaches.
+This performance advantage can be attributed to XGBoost’s inherent strengths in handling heterogeneous tabular data, where its gradient boosting framework effectively captures complex interactions between auxiliary meteorological features, physics-based variables, and high-dimensional image embeddings.
+The ensemble nature of XGBoost enables it to leverage the complementary information from these diverse feature modalities more effectively than traditional neural architectures in this specific nowcasting context.
+Regressor nMAP (%)
+MLP 19.74
+KNN Regressor 16.51
+Random Forest Regressor 15.66
+XGBoost 9.04
+Table 7: Comparison of different regressors for nowcasting using the same input features.
+Results are reported as test nMAP (%, lower is better).
+XGBoost performs the best, indicating its suitability for this task.
+The models are trained on TSI 2020 and evaluated on TSI 2021.
+Appendix B Dataset Details
 Attribute TSI880 Dataset ASI16 Dataset SKIPP’D Dataset
@@
 Number of Samples / Year 24,948 25,107 121,125
-A.1. Overview of Datasets
+Table 8: A Comparative Overview of the TSI880, ASI16, and SKIPP’D Datasets: Key Attributes Including Geographical Location, Data Provided, Image Resolution, Collection Frequency, and Annual Sample Size
+B.1 Overview of Datasets
 TSI880 Dataset: The TSI880 dataset is collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado.
@@
 Images are captured every minute with a resolution of 64×64 pixels, emphasizing finer temporal granularity at the expense of lower image resolution.
-Figure 4.
-Examples of sunny, partly cloudy, and overcast conditions, captured by different sky cameras, are shown from left to right, across the three datasets: TSI, ASI, and SKIPP’D, displayed from top to bottom.
-A.2. Temporal Consistency in Forecasting
+B.2 Robustness to Low-Quality Visual Inputs
+The SKIPP’D dataset helps us assess the model’s robustness under challenging visual conditions.
+With a resolution of just 64×64 pixels, the SKIPP’D images are significantly lower in quality than those found in most practical deployments, where sky imagers typically produce higher-resolution outputs.
+By evaluating on SKIPP’D, we test whether our model can still perform reliably in scenarios where image quality is degraded due to hardware limitations, compression, or transmission artifacts.
+This setup provides a realistic stress test, ensuring that our model is not only accurate under ideal conditions, but also resilient in more constrained, real-world environments.
+Figure 4: Examples of sunny, partly cloudy, and overcast conditions, captured by different sky cameras, are shown from left to right, across the three datasets: TSI, ASI, and SKIPP’D, displayed from top to bottom.
+B.3 Temporal Consistency in Forecasting
 Valid samples for forecasting are formed such that all the data points from time steps $1$ to $T$, and their corresponding forecast intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$, fall within the same day.
@@
 Any samples that violate this condition are considered invalid and are excluded from training or evaluation.
-Appendix B Clear Sky Global Horizontal Irradiance
+Appendix C Clear Sky Global Horizontal Irradiance
 Clear Sky Global Horizontal Irradiance (GHI) is the solar irradiance received on a horizontal surface under cloud-free conditions.
@@
 This becomes very useful in solar forecasting, as this value would give a reference of how much the prediction needs to be.
-Clear Sky GHI is computed using mathematical models incorporating solar position, atmospheric transmittance, and radiative transfer principles.
-A common approach is the Ineichen-Perez model (Stein et al., 2012):
-(5) $GHI_{\text{clear}}=I_{0}\cdot\tau\cdot\cos(\theta_{z})$
+Clear Sky GHI is computed using mathematical models incorporating solar position, atmospheric transmittance, and radiative transfer principles. A common approach is the Ineichen-Perez model :
+$GHI_{\text{clear}}=I_{0}\cdot\tau\cdot\cos(\theta_{z})$ (3)
 where $I_{0}$ is the extraterrestrial irradiance (W/m²), $\tau$ is the atmospheric transmittance factor, $\theta_{z}$ is the solar zenith angle.
-B.1. Extraterrestrial Irradiance ($I_{0}$)
-Extraterrestrial irradiance ($I_{0}$) is the solar irradiance just outside Earth’s atmosphere, slightly varying due to Earth’s elliptical orbit around the Sun. It is given by:
-(6) $I_{0}=S_{c}\cdot\left(1+0.033\cos\left(\frac{2\pi n}{365}\right)\right)$
-where $S_{c}=1367$ W/m² (solar constant), $n$ is the day of the year (1 for January 1, 365 for December 31).
-B.2. Atmospheric Transmittance ($\tau$)
-The atmospheric transmittance $\tau$ accounts for the attenuation of solar radiation by the atmosphere.
-It is often estimated using empirical models, such as the Ineichen-Perez model (Stein et al., 2012):
-(7) $\tau=a\cdot e^{-b\cdot m}$
-where $a,b$ are empirical coefficients dependent on location and aerosol content, $m$ is the air mass, given by (Ineichen and Perez, 2002):
-(8) $m=\frac{1}{\cos(\theta_{z})+0.15(93.885-\theta_{z})^{-1.253}}$
-where $\theta_{z}$ is the solar zenith angle.
-Appendix C Fine-tuning Detailed Results
-C.1. Nowcasting
+Appendix D Fine-tuning Detailed Results
+D.1 Nowcasting
 To understand the impact of fine-tuning duration and the training size, we conducted a series of experiments by varying the amount of training data used for fine-tuning, by using subsets of the data consisting of 1, 2, 3, and 4 weeks.
@@
 Furthermore, in all experimental configurations, our model significantly outperforms the baseline.
-Detailed results for these experiments are presented in Tables 6-9.
-Table 6. Nowcasting Performance with 1 week training
-Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
+Detailed results for these experiments are presented in Tables 9, 10, 11, and 12.
+Trained on Finetuned on SPIRIT wacv2022
 TSI ASI 20.23 52.01
@@
 SKIPP’D 27.51 40.92
-Table 7. Nowcasting Performance with 2 weeks training
-Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 9: Nowcasting performance with one week of training.
+Trained on Finetuned on SPIRIT wacv2022
 TSI ASI 18.96 51.45
@@
 SKIPP’D 26.41 40.25
-Table 8. Nowcasting Performance with 3 weeks training
-Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 10: Nowcasting Performance with 2 weeks training
+Trained on Finetuned on SPIRIT wacv2022
 TSI ASI 16.52 50.38
@@
 SKIPP’D 25.68 39.89
-Table 9. Nowcasting Performance with 4 weeks training
-Trained on Finetuned on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 11: Nowcasting Performance with 3 weeks training
+Trained on Finetuned on SPIRIT wacv2022
 TSI ASI 15.63 50.01
@@
 SKIPP’D 24.32 39.43
-C.2. Forecasting
-Table 10. Forecasting Performance with 2 weeks of training.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 12: Nowcasting Performance with 4 weeks training
+D.2 Forecasting
+Interval Trained on Tested on SPIRIT wacv2022
 1hr TSI ASI 31.15 33.86
@@
 SKIPP’D 36.84 44.16
-Table 11. Forecasting Performance with 4 weeks of training.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 13: Forecasting Performance with 2 weeks of training.
+Interval Trained on Tested on SPIRIT wacv2022
 1hr TSI ASI 22.17 29.03
@@
 SKIPP’D 36.67 44.16
-Table 12. Forecasting Performance with 8 weeks of training.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 14: Forecasting Performance with 4 weeks of training.
+Interval Trained on Tested on SPIRIT wacv2022
 1hr TSI ASI 22.62 32.45
@@
 SKIPP’D 36.23 44.25
-Table 13. Forecasting Performance with 12 weeks of training.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 15: Forecasting Performance with 8 weeks of training.
+Interval Trained on Tested on SPIRIT wacv2022
 1hr TSI ASI 22.03 34.63
@@
 SKIPP’D 36.16 45.28
-Table 14. Forecasting Performance with 16 weeks of training.
-Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
+Table 16: Forecasting Performance with 12 weeks of training.
+Interval Trained on Tested on SPIRIT wacv2022
 1hr TSI ASI 22.76 28.97
@@
 SKIPP’D 38.25 42.46
+Table 17: Forecasting Performance with 16 weeks of training.
 We conducted a series of experiments to assess the impact of training data size on model performance during fine-tuning.
@@
 For each training duration, we performed experiments with different random splits of the corresponding number of weeks and reported the results accordingly.
-The results are presented in Tables 10, 11, 12, 13, and 14.
-Figure 3 was constructed by systematically aggregating the results from our fine-tuning experiments, encapsulating the performance trends observed across different training durations.
+The results are presented in Tables 13, 14, 15, 16, and 17.
+The figure for forecasting performance upon finetuning (Figure 3 in the main paper) was constructed by systematically aggregating the results from our fine-tuning experiments, encapsulating the performance trends observed across different training durations.
 By leveraging visualization techniques, the figure provides a holistic representation of how the model adapts as more site-specific data becomes available.
-It effectively summarizes variations in performance across different random splits of training data and across different sets of source and target datasets.
+It effectively summarizes variations in performance across different random splits of finetuning data and across different sets of source and target datasets.
+This covers training on TSI, and testing on ASI and SKIPP’D, as well as training on ASI, and testing on TSI and SKIPP’D.
+This comprehensive validation methodology substantiates SPIRIT’s efficacy across diverse deployment conditions during transfer learning.
 We employed 95% confidence intervals for all experiments, spanning diverse transfer learning settings and random sampling of the fine-tuning data.
@@
 In every instance, the observed p-values fell below this threshold, demonstrating that SPIRIT achieves statistically significant performance improvements over the baseline.
+D.3 Frozen Vision Encoder in Fine-tuning Strategy
+We perform selective fine-tuning on the regressor and time-series transformer while keeping the vision encoder frozen.
+This is a deliberate methodological choice made to validate our core hypothesis that generalized vision encoders outperform site-specific ones used in prior work.
+Unfreezing the encoder during fine-tuning would introduce location-specific biases and domain-specific adaptations, which would contradict our fundamental objective of demonstrating the superiority of foundation model-based visual representations for cross-domain solar irradiance prediction.
+By maintaining the frozen encoder, we ensure that the observed performance improvements are attributable to the inherent generalizability of pre-trained vision features rather than location and camera configuration specific adaptations, thereby providing a rigorous evaluation of our approach’s transferability across diverse deployment scenarios.
+Appendix E Zero-Shot Transfer Learning Results for More Foundation Models
+To comprehensively evaluate the generalization capabilities of our proposed system, we extend our zero-shot transfer learning analysis beyond the Google ViT Large model to encompass a diverse range of foundation models, including multiple variants of the DINO family (spanning from 22.1M to 1.1B parameters), ConvNeXtV2, and EfficientNet architectures.
+As demonstrated in Table 18, our findings reveal a consistent and compelling pattern across all evaluated models and transfer scenarios.
+While our system demonstrates comparable or incrementally superior performance to the baseline within the same geographical location with the same camera setup, the true strength of our approach emerges when evaluated across different deployment contexts.
+Remarkably, every variation of our system, regardless of the underlying vision encoder architecture, consistently outperforms the baseline method when transferring to new locations and camera configurations.
+This consistent superiority manifests across both cross-camera transfer scenarios, where models trained on TSI2020 are evaluated on ASI2021, and cross-location with cross-task generalization scenarios, exemplified by the transfer from TSI2020 to SKIPP’D2017.
+The robustness of this performance advantage across architecturally diverse models spanning traditional CNN-based approaches like ConvNeXtV2 and EfficientNet, as well as various scales of vision transformer architectures provides compelling evidence for our central hypothesis that generalized vision encoders possess superior transfer learning capabilities for atmospheric forecasting applications.
+This systematic outperformance across different encoder types, parameter scales, and training paradigms underscores the fundamental strength of our system design in leveraging pre-trained visual representations for cross-domain atmospheric prediction tasks.
+The consistency of these improvements across diverse architectural choices validates that our enhanced system design, rather than specific model selection, drives the observed generalization benefits.
+Test Dataset Model 1-Hour 2-Hour 3-Hour 4-Hour
+TSI wacv2022 19.96 22.64 26.30 31.58
+DINO-304M 19.32 21.74 25.29 30.04
+DINO-86.6M 19.52 21.95 25.00 29.57
+DINO-22.1M 20.08 22.53 25.69 29.86
+DINO-1.1B 21.22 23.56 27.93 33.13
+ConvNeXtV2 19.73 22.09 25.17 29.56
+EfficientNet 25.20 27.25 29.58 34.27
+ASI wacv2022 35.74 37.60 37.77 39.58
+DINO-304M 30.11 31.86 34.52 37.91
+DINO-86.6M 30.30 32.01 34.49 38.21
+DINO-22.1M 31.22 33.16 35.60 39.06
+DINO-1.1B 32.41 34.93 36.55 40.12
+ConvNeXtV2 30.42 32.19 34.84 38.91
+EfficientNet 35.06 37.15 39.11 41.94
+SKIPP’D wacv2022 38.88 43.81 47.36 51.78
+DINO-304M 33.20 30.95 32.62 35.71
+DINO-86.6M 33.15 30.79 32.89 35.97
+DINO-22.1M 34.05 31.37 33.21 37.11
+DINO-1.1B 34.95 32.75 34.50 38.78
+ConvNeXtV2 33.71 31.16 32.97 36.12
+EfficientNet 36.51 34.67 35.98 40.29
+Table 18: Zero-Shot Forecasting Results for Multiple Vision Foundation Models (used as encoders in our pipeline) trained on TSI2020 and tested on TSI2021 (TSI), ASI2021 (ASI), and SKIPP’D2017 (SKIPP’D).
+All values represent forecasting error (lower is better).
+Results for the baseline model by wacv2022 are also included for comparison.
+Bold indicates the best value per column within each test dataset.
+Appendix F Implementation Details
+Our implementation incorporates azimuth and zenith angles as auxiliary inputs, from which we derive the physics-based features utilizing the pvlib library: clear sky global horizontal irradiance (GHI), clear sky diffuse horizontal irradiance (DHI), and clear sky direct normal irradiance (DNI) values.
+We then employ these values to obtain the effective irradiance on the panel.
+The auxiliary and physics-based parameters serve dual functions: they generate future covariates and also provide statistical context to the time series encoder regarding past timestamps.
+We use clear sky GHI values as residual inputs to the final MLP layer which contributes favorably to the model’s overall predictive capabilities.
+Comprehensive hyperparameter optimization was conducted across all reported experimental configurations to ensure methodological consistency and reproducibility of results.
+F.1 Hyperparameter Configuration
+For nowcasting, we use XGBoost on a 1297-dimensional feature vector (comprising 1280 ViT features and 17 physics/auxiliary features).
+The hyperparameters are: max_depth = 7, learning_rate = 0.021, n_estimators = 1386, subsample = 0.653, colsample_bytree = 0.888, gamma = 0.002, lambda = 1.744, and early_stopping_rounds = 200.
+For forecasting, we use a 6-layer Transformer with 8 attention heads and a hidden dimension of 1104, along with learnable positional encodings.
+The regressor consists of 10 MLP blocks with residual connections.
+Clear-sky GHI is added as a residual input before the final layer.
+The model is trained using stochastic gradient descent (SGD) with a learning rate of 0.00032, momentum of 0.9, and a warm-up followed by a cosine annealing learning rate schedule.
+Hyperparameter tuning is performed using Optuna.
+F.2 Normalization
+We perform z-score normalization on both the global horizontal irradiance (GHI) and photovoltaic (PV) output signals, standardizing them to have zero mean and unit variance.
+This normalization facilitates stable training and allows the model to directly learn the underlying relationship between sky conditions and energy output without requiring additional bias correction.
+To ensure continuity in time-series inputs, missing image frames or sensor readings are handled by interpolating via the average of their nearest neighboring values, which preserves temporal coherence and minimizes disruptions during model training or inference.
+F.3 Real-Time Deployment Efficiency and Performance
+SPIRIT demonstrates excellent suitability for real-world deployment, with an average inference time of just 0.5378 seconds (±0.0046) on a GPU with 24GB VRAM and 16 CPU cores.
+This low latency ensures that predictions are generated well within the available time window between consecutive image frames.
+In our evaluation datasets, SKIPP’D has the highest temporal resolution with a 1-minute gap between frames, while TSI and ASI, which are representative of most operational sky image datasets, feature 10-minute intervals.
+Given this context, our model’s inference speed is significantly faster than required, confirming that latency is not a bottleneck for real-time solar forecasting applications.
+Appendix G Other Baselines
+G.1 Limitations of Statistical Models
+Classical time series forecasting methods such as ARIMA and Vector AutoRegression (VAR) are frequently used for structured prediction tasks due to their simplicity and interpretability.
+However, these techniques operate solely on temporal sequences and are fundamentally limited in their ability to capture the spatiotemporal complexity inherent in sky imagery.
+These models lack the visual context necessary to model phenomena such as cloud occlusion, movement, and changes, which are critical for accurate solar irradiance forecasting.
+G.2 ARIMA: Univariate Modeling of GHI
+The ARIMA model combines autoregressive (AR) and moving average (MA) components with differencing to handle non-stationary time series.
+For a time series $y_{t}$, the ARIMA(p,d,q)(p,d,q) model is expressed as $\phi(L)(1-L)^{d}y_{t}=\theta(L)\epsilon_{t}$, where $\phi(L)=1-\phi_{1}L-\phi_{2}L^{2}-\cdots-\phi_{p}L^{p}$ is the autoregressive polynomial, $\theta(L)=1+\theta_{1}L+\theta_{2}L^{2}+\cdots+\theta_{q}L^{q}$ is the moving average polynomial, $L$ is the lag operator, and $\epsilon_{t}$ is white noise.
+For our ARIMA(2,0,2)(2,0,2) model applied to GHI, this becomes $y_{t}=\phi_{1}y_{t-1}+\phi_{2}y_{t-2}+\epsilon_{t}+\theta_{1}\epsilon_{t-1}+\theta_{2}\epsilon_{t-2}$, where $y_{t}$ represents the GHI value at time $t$.
+As a statistical baseline, we employ the Autoregressive Integrated Moving Average (ARIMA) model on a univariate time series of Global Horizontal Irradiance (GHI).
+The order $(2,0,2)$ was selected based on Akaike Information Criterion (AIC) and confirmed to be appropriate after verifying stationarity using the Augmented Dickey-Fuller (ADF) test.
+The differencing order $d=0$ was chosen since the time series was found to be stationary.
+The ARIMA model was trained on the TSI 2020 dataset and evaluated on both the TSI 2021 and ASI 2021 datasets.
+The forecasting horizon was set to 4 hours, and only the preceding 1 hour of data (equivalent to 6 frames with 10-minute intervals) was used as context, in alignment with SPIRIT’s input and output windows.
+The ARIMA model yielded a Mean Absolute Error (MAE) of 168.94, a Root Mean Square Error (RMSE) of 195.13, and a negative $R^{2}$ score of -0.6180, with a normalized Mean Absolute Percentage (nMAP) of 55.51%.
+These results were consistent across both the TSI 2021 and ASI 2021 datasets.
+This is expected, as the ARIMA model operates purely on GHI values and does not utilize any image-based input.
+Since the datasets share the same auxiliary irradiance data and originate from the same NREL facility in Colorado, the only difference in camera configuration has no effect on this statistical model’s performance.
+G.3 VAR: Multivariate Time Series Forecasting
+The Vector AutoRegression (VAR) model extends univariate autoregressive modeling to multivariate time series, where each variable is regressed on its own lagged values and the lagged values of all other variables in the system.
+For a VAR(p)(p) model with $k$ variables, the system is represented as $\mathbf{y}_{t}=\mathbf{c}+\mathbf{A}_{1}\mathbf{y}_{t-1}+\mathbf{A}_{2}\mathbf{y}_{t-2}+\cdots+\mathbf{A}_{p}\mathbf{y}_{t-p}+\mathbf{e}_{t}$, where $\mathbf{y}_{t}$ is a $k\times 1$ vector of endogenous variables at time $t$, $\mathbf{c}$ is a $k\times 1$ vector of constants, $\mathbf{A}_{i}$ are $k\times k$ coefficient matrices for $i=1,2,\ldots,p$, and $\mathbf{e}_{t}$ is a $k\times 1$ vector of white noise error terms.
+In our implementation, $\mathbf{y}_{t}$ contains clear-sky GHI, clear-sky DNI, actual GHI, and physics-based auxiliary variables, enabling the model to capture interdependencies between these solar irradiance components.
+We also evaluate a Vector AutoRegression (VAR) model trained on multivariate time series composed of clear-sky GHI, clear-sky Direct Normal Irradiance (DNI), actual GHI, and physics-based auxiliary variables.
+To maintain consistency with the image-based baselines, we limit the context window to the previous 6 time steps (1 hour), setting maxlags=6.
+The VAR model achieved an MAE of 147.25, RMSE of 174.99, an $R^{2}$ value of -0.3013, and a normalized MAP of 48.38% on both the TSI 2021 and ASI 2021 datasets.
+This uniformity in performance is due to the shared irradiance and auxiliary variables used as input, which are identical across the two datasets.
+As with ARIMA, the VAR model does not utilize visual data, and therefore its forecasting capability is unaffected by the differing camera configurations of TSI and ASI.
+G.4 The Limitations of Purely Statistical Approaches
+Despite being well-established in traditional forecasting pipelines, autoregressive statistical models, whether univariate (ARIMA) or multivariate (VAR), fall significantly short when tasked with predicting solar irradiance.
+Their inability to process visual inputs makes them fundamentally incapable of modeling the intricate spatial patterns and dynamic cloud formations captured in sky images.
+Both ARIMA and VAR produce high error rates and even negative $R^{2}$ values, highlighting their limited generalization capabilities.
+In contrast, vision-based models like ours exhibit strong predictive accuracy even under location shifts.
+These findings underscore the necessity of incorporating spatiotemporal visual cues for solar irradiance forecasting.
+G.5 On CNN-LSTM Baselines
+We benchmark our approach against the current state-of-the-art method (wacv2022), which has demonstrated superior performance on publicly accessible datasets and established the performance ceiling for solar forecasting tasks.
+Vision transformer architectures have fundamentally superseded CNN-LSTM approaches across forecasting domains, with (wacv2022) which outperforms traditional convolutional methods (talha2019).
+Our experimental design prioritizes meaningful evaluation against the strongest available competitor, ensuring that any performance gains represent genuine technological advancement in solar forecasting capabilities.
+The established dominance of transformer-based architectures in computer vision forecasting tasks makes comparisons with CNN-LSTM baselines redundant as they have been systematically outperformed.
+Our focused comparison allows for a clear analysis of architectural innovations in solar forecasting, without aiming to reestablish existing performance hierarchies.
+Appendix H Detailed Results with Standard Errors
+Table 19 reports the detailed forecasting performance of our model, including standard errors computed across multiple random trials.
+The consistently small standard errors across all forecast horizons demonstrate the robustness and stability of our approach.
+In contrast, Table 20 presents results obtained when the model is trained and evaluated without future covariates, showing markedly degraded accuracy.
+Importantly, the performance gaps between the two settings are substantially larger than the associated standard errors, indicating that the observed improvements are significant.
+This underscores the critical role of future covariates in enhancing predictive accuracy for longer forecast horizons.
+Forecast Horizon Mean ± Std. Error
+Nowcasting 9.080 ± 0.00526
+1-hour forecast 19.136 ± 0.092
+2-hour forecast 21.779 ± 0.101
+3-hour forecast 25.317 ± 0.170
+4-hour forecast 29.780 ± 0.259
+Table 19: Detailed forecasting performance of our model with mean values and standard errors across random trials.
+The results show consistent improvements with low variance for nowcasting and future forecasts.
+Forecast Horizon Mean ± Std. Error
+1-hour forecast 19.688 ± 0.085
+2-hour forecast 22.858 ± 0.088
+3-hour forecast 26.800 ± 0.083
+4-hour forecast 31.890 ± 0.095
+Table 20: Forecasting performance of the model without future covariates, showing degraded accuracy across horizons.
+The gap between these results and the main model highlights the importance of future covariates.
+Appendix I Comprehensive Analysis of Evaluation Metrics
+The evaluation of solar forecasting models requires careful consideration of appropriate metrics that accurately reflect model performance in real-world deployment scenarios.
+We employ multiple evaluation metrics to provide a comprehensive assessment of our proposed methodology.
+The normalized Mean Absolute Percentage error (nMAP) is defined as:
+$\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$ (4)
+Additionally, we evaluate performance using Mean Absolute Error (MAE):
+$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$ (5)
+Root Mean Square Error (RMSE):
+$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$ (6)
+and coefficient of determination ($R^{2}$):
+$R^{2}=1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$ (7)
+where $y_{i}$ represents the actual value, $\hat{y}_{i}$ represents the predicted value for the $i$-th sample, and $\bar{y}$ is the mean of actual values, with $i\in\{1,\ldots,N\}$.
+The comprehensive results across all these metrics are presented in Table 21.
+Interval Tested on nMAP RMSE R2 MAE
+1 hr TSI 2021 18.96 176.60 0.56 106.63
+ASI 2021 29.99 200.05 0.50 129.11
+SKIPP’D 32.93 205.11 0.42 134.19
+2 hr TSI 2021 21.77 195.07 0.49 120.80
+ASI 2021 31.71 203.01 0.43 132.06
+SKIPP’D 29.01 198.53 0.47 127.54
+3 hr TSI 2021 25.46 201.73 0.53 129.66
+ASI 2021 34.41 207.59 0.45 136.87
+SKIPP’D 30.26 200.56 0.52 129.58
+4 hr TSI 2021 29.89 199.73 0.60 128.77
+ASI 2021 38.00 214.55 0.58 143.26
+SKIPP’D 34.63 207.96 0.61 137.21
+Table 21: SPIRIT’s generalization performance across datasets and time intervals.
+The model is trained on TSI 2020 and evaluated on TSI 2021, ASI 2021, and SKIPP’D using four forecasting horizons (1 hr–4 hr).
+We report four metrics: nMAP, RMSE, R2, and MAE.
+We primarily focus on nMAP as our principal evaluation metric due to its established status as the standard metric in solar forecasting research.
+This standardization facilitates meaningful comparisons across different methodologies and ensures our contributions align with established benchmarks in the field.
+The practical significance of nMAP improvements extends far beyond academic metrics, as even marginal improvements translate to substantial economic and environmental benefits when deployed across large-scale solar installations.
+These improvements compound across extensive solar deployments, potentially saving millions in operational costs while simultaneously reducing carbon emissions through enhanced grid stability and reduced reliance on fossil fuel backup systems.
+The normalized nature of nMAP ensures that our improvements maintain consistent significance across diverse geographic conditions and varying solar irradiance patterns.
+This geographic invariance is crucial for developing universally applicable forecasting models that can be deployed across different climatic zones and solar resource conditions without requiring region-specific recalibration of performance expectations.
+While we conduct comprehensive multi-metric evaluation during model development and validation as shown in Table 21, we deliberately emphasize nMAP as the decisive evaluation metric throughout our analysis to maintain consistency with established solar forecasting standards and provide clear, interpretable results.
+Appendix J Other Details
+Longer forecasts: SPIRIT supports longer-horizon forecasts.
+However, our problem setting focuses on short-term forecasting, critical for grid operators who need high-resolution predictions for real-time dispatch decisions and load balancing.
+Beyond 4 hours,the focus shifts away from grid control, and accuracy deteriorates due to atmospheric randomness and clouds moving beyond the camera’s field-of-view, which is why longer-term forecasting relies on numerical weather prediction(NWP) and satellites, representing a fundamentally different forecasting paradigm.
+Zero-Shot Transfer and Historical Data Requirements:
+To clarify our methodology’s data requirements, our approach eliminates the need for historical data at new deployment sites.
+New solar installations typically require 5-10 years of site-specific data before providing reliable forecasts, creating a major bottleneck in renewable energy expansion and delaying the deployment of effective forecasting systems.
+SPIRIT addresses this fundamental challenge by leveraging datasets from established sites to achieve zero-shot transfer to new locations, enabling immediate deployment of forecasting systems without any data collection period.
+This capability is crucial for accelerating solar energy adoption, as it eliminates the traditional waiting period required for accumulating sufficient historical data at each new installation site.
+While our model does utilize historical data during training from established sites, the key innovation lies in transferring this learned knowledge to new sites without requiring any site-specific historical data collection.